### v3

# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

Abstract
Warning: This paper includes examples that may be deemed sensitive or offensive.
Aligning language models (LMs) with human opinion is challenging yet vital to enhance their grasp of human values, preferences, and beliefs. We present ChOiRe [^1] Codes and data will be available at https://github.com/dxlong2000/ChOiRe., a four-step framework to predict human opinion which differentially models the user’s explicit personae (i.e. demographic or ideological attributes) that are manually declared, and implicit personae inferred from user historical opinions. ChOiRe consists of (i) an LM analyzing the user’s explicit personae to filter out irrelevant attributes; (ii) the LM ranking the implicit persona opinions into a preferential list; (iii) Chain-of-Opinion (CoO) reasoning, where the LM sequentially analyzes the explicit personae and the most relevant implicit personae to perform opinion prediction; (iv) and where ChOiRe executes Step (iii)’s CoO multiple times with increasingly larger lists of implicit personae to overcome insufficient personae information to infer a final result. ChOiRe achieves new state-of-the-art effectiveness with limited inference calls, improving previous techniques significantly by 3.22%. We also show that ChOiRe’s Steps (i) and (ii) can significantly better fine-tune opinion-aligned models, by up to 18.44%.
1 Introduction
Language models (LMs) are becoming indispensable tools, serving in various roles such as dialogue agents OpenAI (2022); Google (2022), data analysts Wang et al. (2023a); Cheng et al. (2023), and decision support Ye et al. (2023). LMs also demonstrate the capability to model distinct opinions which influence response generation on input queries Bai et al. (2022); Glaese et al. (2022); Santurkar et al. (2023). Unfortunately, the opinions modeled by language models are shaped by the extensive training and feedback data, which are themselves influenced by countless human perspectives, making them inherently challenging to model. As human–AI interactions become common, it becomes imperative to align models with human opinion to meet individual and personalized expectations.
Alignment frameworks such as RLHF Christiano et al. (2017); Ouyang et al. (2022) have been devised to train and fine-tune LMs. Nonetheless, applying them to align large language models (LLMs) with human opinions is difficult because they require both significant computational resources and high-quality supervised feedback data, which is itself challenging to collect. Therefore, prompt-based opinion alignment through persona has been studied as an easy-to-deploy and easy-to-combine resource-parsimonious alternative Perez et al. (2023); Simmons (2023); Santurkar et al. (2023); Deshpande et al. (2023).
However, even when aligning with well-represented groups, persona-based methods result in low steerability Santurkar et al. (2023). This raises significant concerns when using LMs to model for individual users. Since individuals hold nuanced opinions that evolve and are influenced by situational factors, it is more challenging to achieve individual opinion LM alignment. Hwang et al. (2023) find significant opinion variations among individuals sharing the same demographics, exposing flaws in current group-focused alignment. They argue for individualized models, introducing an approach (detailed below) integrating user’s demographic and ideological attributes, (we term as explicit personae) and historical opinions (implicit personae) into the prompt for opinion prediction.
While Hwang et al. (2023)’s naïve integration achieves promising results, we argue that it suffers from a few key limitations. First, it employs all explicit personae. However, we contend that only a subset is needed for accurate opinion prediction; including non-relevant personae may act as noise, harming predictive performance. Second, Hwang et al. (2023) utilize the top-$K$ semantically similar opinions with the question (here termed top-$K$ implicit personae). This approach is inefficient, as such similar opinions may not offer the most valuable information for opinion prediction. Additionally, our empirical experiments suggest that LMs may lack sufficient personae evidence with a fixed $K$ — dynamically adjusting $K$ per task can overcome such deficiencies. Finally, while Chain-of-Thought (CoT; Wei et al. 2022; Kojima et al. 2022) enables LMs to perform multi-step reasoning tasks effectively, we find that the naïve application of CoT does not help modern LLMs like ChatGPT with opinion alignment [^2] We validate all our claims on these limitations in Sections E.1, 6.1, 4 and 5, respectively..
To address the above challenges, we propose ChOiRe [^3] Chain of Opinion Reasoning, pronounced as the English word “choir”.(fig. 1), a four-step solution for opinion prediction leveraging LLMs’ strong data evaluation and analytic capabilities Wang et al. (2023a); Cheng et al. (2023). First, an LLM analyzes a target user’s explicit personae to discard irrelevant ones. Second, the LLM ranks implicit persona opinions in order of usefulness, selecting the top-$K$ as the most valuable. This surpasses the constraint of using semantic similarity scores. Third, we introduce Chain-of-Opinion (CoO), a designed variant of CoT that allows the LLM to explain and analyze selected explicit personae and top-$K$ implicit personae sequentially. ChOiRe applies self-consistency over CoO to provision the appropriate amount of user information for opinion inference.
ChOiRe achieves new state-of-the-art (SOTA) in opinion alignment effectiveness and reliability, while using a limited inference budget. We conduct a thorough analysis to verify our hypotheses concerning explicit and implicit personae and defend our Chain-of-Opinion reasoning methodology. Our analysis further reveals that ChOiRe’s first two steps in handling explicit and implicit personae also helps to fine-tune opinion-aligned models.
2 Related Work
Aligning LMs with Humans.
Aligning language models with human behaviour is a recent area of study as alignment can increase user experience satisfaction and utility Wang et al. (2023c). One line of work develops prompting techniques with user demographic information (e.g., political identity) to encourage LMs to output human-like responses. Argyle et al. (2023) show that by properly conditioning LMs with targeted identity and personality profiles, it is possible to produce biased outputs that strongly correlate with human responses. Furthermore, Simmons (2023) claims that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases. Despite recent advances, Santurkar et al. (2023) discovered that LMs align poorly with human opinions, as evidenced by model performance on public opinion polls. Hwang et al. (2023) recently propose to incorporate explicit and implicit personae to predict human opinions in new contexts. In section 1, we argue that this naïve strategy is suboptimal. ChOiRe overcomes these limitations.
Reasoning with LMs via Prompting.
Large-scale model architectures Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023) have enabled large language models (LLMs) to excel at various NLP tasks using zero- or few-shot prompting Liu et al. (2023). Notably, Wei et al. (2022); Kojima et al. (2022) propose prominent Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps to solve multi-step reasoning tasks with higher fidelity and efficiency.
Can CoT analyze and predict human opinion effectively? We find that a naïve application of CoT does not help GPT-X models (section 5), but that an appropriate modification does. We propose Chain-of-Opinion (CoO) reasoning (section 3) that overcomes CoT’s limitations in this task. Noting that prompting techniques such as task decomposition Khot et al. (2023); Zhou et al. (2023) and retrieved-based methods Yao et al. (2023); Shinn et al. (2023) have been recently introduced, we focus only on the reasoning explanation aspect here given the abstractive and challenging nature of the task.
Figure 1: ChOiRe overview, consisting of the four main steps (cyan background), as detailed in section 3.
3 ChOiRe: A Chain of Opinion Framework
Task Formalisation. We follow Santurkar et al. (2023), and formulate the opinion prediction task as multiple-choice question answering. Formally, a benchmark with $N$ data points is denotated as $D=\{\langle T,E,I,q,a\rangle_{n}\}_{n=1}^{N}$, where $T$, $E$ and $I$ indicate the (T)opic of a question $q$, the (E)xplicit personae and (I)mplicit personae of the user answering $q$ with opinion $a$. Following the prior work, $E$ consists of $12$ user demographic and ideology metadata attributes, and $I$ contains a number of the user’s historical opinions in the format of question–answer pairs. Models then learn to analyze $T,E,I,q$ and predict the opinion $a$.
Fig. 1 shows an overview of ChOiRe, consisting of four main steps (marked with a cyan background). First, ChOiRe employs an LLM to analyze and select a subset of relevant explicit personae, denoted as $E^{rel}\subseteq E$ for answering the opinion question $q$. The LLM then assesses the informativeness of the implicit personae ($I$) in predicting $q$, selecting the top-$K$ implicit personae (termed LLMtop-$K$). Next, an LLM is prompted to explain the provided explicit $E^{rel}$ and implicit LLMtop-$K$ personae sequentially in a Chain-of-Opinion (CoO) reasoning strategy. Finally, ChOiRe calls the LLM to predict the opinion $a$ with varying values of $K$ for the top-$K$ implicit personae. ChOiRe chooses the opinion with the highest frequency as the final prediction We detail the steps below.
3.1 Filtering Explicit Personae Attributes (FEA)
Accounting for explicit personae, which consist of the demographic and ideological metadata attributes of users — such as their age, income, and political ideology — has been shown to help models characterize and predict human opinions more accurately Hwang et al. (2023). However, which personae matter and which do not? are still open questions. Section E.1 shows such an example in full, where, when considering all of explicit personae, the model makes an incorrect prediction while removing unnecessary personae the model made a correct prediction. This may be caused by the LLM’s attention mechanism’s forcing the model to attend to all input tokens, even irrelevant ones. To filter out unnecessary explicit personae, we ask the LLM to reason and analyze how each persona is helpful for the model to predict the opinion via Chain-of-Thought to output a list of which personae are relevant given the question and the opinion answer choices [^4] We provide our FEA prompt in Section C.1. Surprisingly, we find that LLMs evaluate more than half of the explicit personae as irrelevant on average. We further conduct human evaluations to verify this finding in section 4.
3.2 Implicit Personae Opinions Ranking (LLMtop-$K$)
LLMs have been sensitive to selected demonstrations and their order in the prompts Perez et al. (2021); Luo et al. (2023); Gao et al. (2023). For predicting human opinions, we discover that LLMs are also sensitive to the chosen implicit personae opinions as input. Hwang et al. (2023) rank the implicit personae opinions via semantic-similarity scores and select top-$K$. This strategy is suboptimal because the top-ranked opinions in terms of semantic similarity may not be the ones that provide the most supportive information for the models to predict opinions (section E.2). As LLMs are shown to be good data analysts Wang et al. (2023a); Cheng et al. (2023), we propose to address the above challenge by utilizing LLMs to analyze and rank the implicit personae opinions in usefulness descending order. Our finding is that despite the output rankings from LLMs varying with different input orders of implicit personae opinions, the sets of LLMtop-$K$ opinions overlap by a good coefficient when $K$ is large enough ($\geq 8$) (fig. 5). Therefore, we propose to input the implicit personae opinions to LLMs in a random order to make our method more versatile. We also examine the case where we input the opinions in the semantic similarity order. We illustrate the prompt template in section C.2. By performing this step, our proposed method supports the usefulness of opinions in predicting the test opinions, rather than the semantic similarity. We term this method as LLMtop-$K$.
3.3 Chain-of-Opinion Reasoning (CoO)
Wei et al. (2022); Kojima et al. (2022) introduce few-shot and zero-shot Chain-of-Thought (CoT) prompting strategies demonstrating that by reasoning step-by-step, LLMs can achieve promising results on complex tasks. However, the sampled reasoning steps can be inconsistent, leading to possibly different outcomes Wang et al. (2023b). Furthermore, it is little known how the models perceive multiple implicit personae opinions, especially when many opinions are input, which one(s) the models used, which one(s) they didn’t for predicting the opinion? Our preliminary experiments with CoT (sections 6.1, E.4 and E.3) reveal that the CoT explanations can vary frequently based on different subsets of opinions mentioned in their explanations, leading to diverse final answers, especially when the decoding temperature is relatively high555greater or equal to 0.6. To mitigate this issue, we propose to instruct the LLMs to analyze the given explicit and implicit personae one by one before concluding the prediction via simply adding "explaining and analyzing how each of the Opinions and Demographic Information supports the question" into the prompt instruction. Given an LLM that can follow human instructions well such as ChatGPT OpenAI (2022), this addition offers two notable advantages. First, for each question, we ensure that the model explains and analyzes the provided personae one by one without missing any, possibly resulting in more thorough predictions. Second, this method helps the model to output more consistent reasoning explanations, enhancing its reliability (section 6.1).
3.4 Answer Consistency with Dynamic Numbers of Opinions
Prior work Hwang et al. (2023) fixes the number of implicit personae opinions for prediction to $K=8$. However, this approach occasionally results in models generating "...the answer cannot be determined." (tables 4 and E.5). We attribute this to insufficient user implicit personae opinions provided. Inspired by Self-Consistency (SC) Wang et al. (2023b), our approach involves sampling multiple answers using different $K$ values for a given question. The most frequent answer, along with its explanation, becomes the final prediction. Our method is distinct from SC since SC samples multiple answers with a fixed prompt. We experiment with $K\in\{8,10,12\}$ for efficiency.
4 Evaluation
Dataset.
We experiment on OpinionQA dataset Santurkar et al. (2023) — the only opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups. It covers $60$ US demographic groups, with $15$ topics, each comprising around $100$ questions, gathered from $5,340$ users.
Dataset Preprocessing.
Due to limited resources, we randomly sample $25$ users per topic for our experiments. For each user, we follow Hwang et al. (2023) to use $20\%$ of the implicit questions as the implicit persona. For the remaining $80\%$ implicit questions, we randomly select a maximum of $15$ implicit questions for testing. Our sampling method results in a total of $375$ users and $5,603$ implicit evaluation question–answer pairs. Our subset is highly representative because we gather 25 users from every topic and 15 questions per user. Rigorous statistical tests further validate the significance of our results which align closely with Hwang et al. (2023) testing on a larger subset using InstructGPT.
Prompting Baselines.
We use both closed-source ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b), and open-source Mistral-7B-Instruct-v.02 Jiang et al. (2023) as our LLMs, and compare ChOiRe with $5$ prompting methods: (1) W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; (2) Demographic + Ideology + top8 Opinions (termed DIO-top8), introduced by Hwang et al. (2023) demonstrating that integrating explicit and implicit personae enhances user opinion modeling and prediction, achieving state-of-the-art results on OpinionQA at that time; (3) DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al. (2022) version of DIO-top8 involving appending "answer the following question step-by-step" to prompts, aiming to explore whether CoT improves model performance in this task; (4) DIO-top8 + SC is the baseline which we apply the Self-Consistency technique with CoT Wang et al. (2023b) to DIO-top8 to select the most frequent answer generated by the model as the final opinion prediction; (5) DIO-top8 + Self-refine Madaan et al. (2023) interactively feedbacks and refines the answers by LLMs. We do not experiment with InstructGPT Ouyang et al. (2022) like Hwang et al. (2023) since this model is going to be deprecated and replaced by ChatGPT-Instruct. For GPT-4, we only run the main experiment and we use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget. All the prompts and costs are in appendix C, implementation details in section A.1, and more baselines in section B.1.
Fine-tuning Baselines.
We further investigate whether ChOiRe’s FEA and LLMtop-$K$ steps (section 3) also improve fine-tuning for opinion-aligned models. We first create the fine-tuning data by using ChatGPT to perform ChOiRe’s FEA and LLMtop-$K$ steps ($K=8$) on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones. We then fine-tune and evaluate GPT-2 models (base, large) Radford et al. (2019) and FlanT5 models (base, large) Chung et al. (2022). Fine-tuning details are provided in section A.1.
Metrics.
We employ Accuracy and Collapsed Accuracy as the automatic evaluation metrics following Hwang et al. (2023). It is worth noting that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples. In addition, human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$ and CoO steps of ChOiRe. Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via CoO. To this end, we randomly select $100$ answers generated by ChOiRe with ChatGPT, ChatGPT-Instruct, GPT-4, Mistral. We then hire $3$ excellent undergraduates who are native English speakers as annotators. For FEA and LLMtop-$K$ steps, each annotator is instructed to rate on a 1-3 scale (3 is best) via the Satisfaction criterion defined as how well the algorithm of LLMs performs in filtering/ranking, subjectively. To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the CoO explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions. Three annotators are also guided to rate the criteria on a 1-3 scale. Each metric’s final score is the average of three annotators’ scores. The scoring instructions are in section D.1 and the inter-annotators’ agreement is assessed by Kripp’s alpha Krippendorff (2011).
5 Main Results
Model ChatGPT ChatGPT-Inst GPT-4 Mistral-7B-Ins.-v0.2
W/o persona 46.60/65.72 44.91/63.60 - 41.24/59.54
DIO-top8 50.22/69.21 51.95/71.16 57.98/76.86 44.16/62.47
DIO-top8 + Self-refine 43.14/65.33 42.71/62.98 - 36.23/55.06
DIO-top8 + CoT 49.96/69.05 51.90/71.51 - 52.25/71.95
DIO-top8 + SC 50.58/69.66 52.06/71.87 - 53.14/72.88
DIO-top8 + FEA 50.64/69.85 52.63/72.30 - 44.99/64.09
DIO-top8 + CoO 50.97/70.22 52.08/71.65 - 53.79/73.59
DIO-LLMtop8 51.03/70.31 52.80/72.60 - 45.86/64.98
DIO-LLMtop8 + FEA 51.19/70.69 52.97/72.84 - 45.23/64.73
DIO-LLMtop8 + FEA + CoO 51.90/71.57 53.01/72.91 59.02/78.70 54.21/74.09
ChOiRe 52.21normal-$\dagger$/72.09normal-$\dagger$ 53.26normal-$\dagger$/73.26normal-$\dagger$ 59.30normal-$\dagger$/78.82normal-$\dagger$ 54.43normal-$\dagger$/74.34normal-$\dagger$
% Improvements +3.22/+3.49 +2.52/+1.93 +2.28/+2.55 +2.42/+2.00
Table 1: Overall Accuracy/Collapsed Accuracy on ChatGPT, ChatGPT-Instruct, and GPT-4. FEA is our first step, Filtering Explicit Attributes. LLMtop8 is the second step, ranking and selecting top-8 implicit persona opinions as demonstrations, CoO is Chain-of-Opinion reasoning. Improvements are calculated with the best baseline. $\dagger$ denotes our model outperforms baselines significantly with p-value < 0.01 under t-test.
Model GPT-2-base GPT-2-large FlanT5-base FlanT5-large
W/o persona 41.14/58.87 21.94/39.11 48.98/68.33 39.83/58.43
DIO-top8 21.23/38.64 24.94/42.22 55.00/74.98 54.94/74.79
DIO-top8 + FEA 22.62/40.97 25.65/45.21 55.78/75.34 58.77/77.26
DIO-LLMtop8 22.65/41.12 28.86/47.60 57.97normal-$\dagger$/77.46normal-$\dagger$ 58.20/77.56
DIO-LLMtop8 + FEA 25.05/44.41 29.54normal-$\dagger$/48.66normal-$\dagger$ 57.45/77.13 59.00normal-$\dagger$/78.46normal-$\dagger$
% Imp. over DIO-top8 +17.99/+14.93 +18.44/+15.25 +5.40/+3.30 +7.38/+4.90
Table 2: Performance of fine-tuned baselines with our proposed FEA and LLMtop8 steps preprocessed by ChatGPT. $\dagger$ denotes our model significantly outperforms baselines with p-value < 0.01 under t-test.
Overall Prompting Results.
Table 1 shows our macro experimental outcomes. We derive $4$ main observations in this task. First, ChOiRe improves the best among baselines significantly with $3.22\%$, $2.52\%$, $2.28\%$, $2.42\%$ accuracy for ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. It establishes a strong SOTA result with GPT-4, surpassing previous SOTA DIO-top8 with InstructGPT achieving 53.74% Hwang et al. (2023) by a notable margin. Notably, in the case of GPT-4, we utilize ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one. Second, we see that Accuracy and Collapsed Accuracy have the same trend, and ChOiRe also achieves the SOTA on Collapsed Accuracy with the highest improvement of $3.94\%$ observed with ChatGPT. Third, naïve CoT ("answer the following question step-by-step") helps Mistral but slightly harms ChatGPT and ChatGPT-Instruct with DIO-top8 Hwang et al. (2023). On the other hand, SC improves all models. Therefore, we attribute CoT’s limitation to the inconsistency of its explanations (3). Meanwhile, ChOiRe with CoO consistently attains improvements, verifying the effectiveness of explicitly requiring the model to analyze all the personae. Finally, ChatGPT, ChatGPT-Instruct, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average, respectively. This suggests that over half of explicit personae may be noisy for models to predict opinions.
Fine-grained Prompting Results.
Diving deeper into the benchmark topics in table 5, ChOiRe achieves SOTA results in $8/15$, $8/15$, $11/15$, $13/15$ topics for ChatGPT and ChatGPT-Instruct, GPT-4, and Mistral. The improvements are especially huge for some topics. For example, compared with the best among baselines, it improves GPT-4 up to $12.08\%$ accuracy on Views on gender, ChatGPT up to $9.82\%$ on Economic Inequality. We also specifically compare ChOiRe with the best baseline DIO-top8 + SC in section B.3, showing $8/12$ improvements for ChatGPT and ChatGPT-Instruct. We further plot the accuracy distribution over users of ChOiRe, specifically for ChatGPT in fig. 4. We see that the majority accuracy is $0.5$, with a few users scoring zero and over $20$ achieving perfection.
Fine-tuning Results.
Table 2 presents our fine-tuning outcomes. Notably, leveraging the ChOiRe’s FEA and LLMtop-$K$ steps on the fine-tuning data yields substantial enhancements for GPT-2-large and FlanT5-large, showcasing relative accuracy improvements of $18.44\%$ and $7.38\%$ respectively. Remarkably, ChOiRe’s FEA and LLMtop-$K$ steps bring FlanT5-large’s performance on par with GPT-4, despite GPT-4’s significantly stronger capability. Furthermore, ChOiRe’s LLMtop-$K$ proves particularly beneficial for enhancing FlanT5-base. Surprisingly, GPT-2-base performs well even without user demographic and ideological information, possibly due to potential contamination Sainz et al. (2023) with public polling data from OpinionQA.
Human Evaluation Results.
Model FEA Satis. LLMtopK Satis. Rea. Foll. Inst.
[…]
Our human evaluation results in table 3 reveal three key findings. First, ChatGPT and ChatGPT-Instruct achieve similar performance in filtering explicit personae and ranking opinions, while Mistral achieves lower results. While ChatGPT excels slightly in ranking, ChatGPT-Instruct performs slightly better in explicit personae selection. Three models proficiently filter unnecessary explicit personae, but ranking opinions poses a more challenging task intuitively and empirically, with a common error being the inconsistent relevance ranking of opinions, sometimes misplacing high-level relevance. Second, four models effectively generate reasonable thoughts leading to the final answer, and GPT-4 performs the best. Finally, ChatGPT and ChatGPT-Instruct follow our instructions to explain and analyze the explicit and implicit personae provided one by one with CoO significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$. We hypothesize that this is because ChatGPT and ChatGPT-Instruct excel in following instructions, while GPT-4 is optimized for completing texts.
6 Discussion
We discuss the main analyses in this section. Extra analyses are presented in appendix B.
6.1 Methodology Analysis
Ablation of FEA.
To gauge the impact of filtering unnecessary explicit personae (FEA) on performance, we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al. (2023), denoted as DIO-top8 + FEA in table 1. The results indicate enhancements with DIO-top8 + FEA achieving a 0.8%, 1.3%, 1.9% accuracy performance boost on ChatGPT, ChatGPT-Instruct, and Mistral respectively. This underscores the effectiveness of eliminating irrelevant explicit personae in improving the models’ ability to understand and predict human opinions.
FEA via Topics.
To understand the explicit personae filtered by LLMs across various topics, we document the top-3 removed personae in section B.5. We observe that "Citizenship" is consistently the most frequently removed attribute, followed by "Race". This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation. Another explanation may be the lack of correlation between citizenship/race and opinions in the US-centric OpinionQA dataset. Additionally, we also see that ChatGPT often categorizes ‘‘Marital status" as non-useful, ChatGPT-Instruct commonly removes ‘‘Frequency of religious attendance", and ‘‘Gender" got removed by Mistral, revealing potential biases in LLMs.
LLMtop-$K$ versus Top-$K$.
From table 1, DIO-LLMtop8 outperforms DIO-top8 by $1.6\%$, $1.6\%$, $3.8\%$ accuracy on ChatGPT, ChatGPT-Instruct, Mistral confirming that prioritizing meaning and usefulness improves opinion prediction. One possible explanation for this can be the orders ranked by semantic similarity scores only consider ranking with the input questions Hwang et al. (2023), while our orders consider both input questions and their answer choices (fig. 1). We further explore two key aspects: (1) The agreement of LLM-orders and semantic similarity orders, and (2) Points of maximum disagreement between these orders. To measure the ranking agreements, we calculate Kendall’s Tau correlation coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-Instruct, and Mistrial and orders sorted by semantic similarity scores, and the results are presented in fig. 6 and fig. 7. Surprisingly, for ChatGPT and ChatGPT-Instruct, we find that the two ranking orders have minimal monotonous relations with means approximating 00 and low standard deviations showing no agreement. For Mistral, we find a low agreement with a mean of $0.43$ score. These low and no agreements further verify that ranking by usefulness can be very different from ranking by semantic similarity. We also deep dive into cases with notable order variations to address (2). Section E.2 illustrates one such case in the "Guns" topic. We derive three observations. First, not all top-8 opinions by semantic similarity scores help predict the opinion. For example, the $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
Opinions Order Analysis of LLMtop-$K$ Step.
The performance difference of DIO-top8 and DIO-LLMtop8 in table 1 highlights that LLMs are sensitive to the chosen implicit personae opinions. An important question arises: Are LLMs also affected by the input order of implicit persona opinions in the ranking step (section 3)? Our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large. We randomly select $300$ questions, shuffle implicit persona opinions four times with different seeds, and record four LLM ranking outputs for each. We also collect one more LLM ranking output by feeding implicit personae opinions in semantic similarity order. For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap coefficient Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$. The scores, shown in fig. 5, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models. Despite this, is there substantial variance in model performance across random seeds? Our findings reveal no significant variance, with the variants statistically outperforming the baseline DIO-top8. Specifically, we assess ChatGPT and Mistral with DIO-LLMtop8 on $3$ out of $4$ random seeds, detailed in section B.6. The results demonstrate relatively small standard deviations in their performance, and critical values of $99\%$ CI of DIO-LLMtop8 under t-test for both models surpass DIO-top8, confirming that LLMtop8’s effectiveness is not due to randomness.
CoO versus CoT.
Figure 2: Consistency scores of the baseline DIO-top8 (ChatGPT) with CoO and CoT.
Table 1 indicates that Chain-of-Thought (CoT) Kojima et al. (2022) slightly harms baseline DIO-top8 performance for ChatGPT and ChatGPT-Instruct. Conversely, our Chain-of-Opinion reasoning (CoO) enhances overall performance for all models. To investigate the consistency of CoT and CoO, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and CoO, at $3$ different temperatures $0.3$, $0.6$, $0.9$. For each prompting technique, we measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score. The results are illustrated in fig. 2 showing that CoO brings slightly better consistent answers compared to CoT, especially when the temperature is high which verifies CoO potentially enhances the reliability of LLMs.
Dynamic Numbers of Opinions Analysis.
Table 4 illustrates our analysis answering two research questions: (1) How frequent can’t LLMs answer the question? and (2) How do LLMs perform when more opinions than $K=8$ are provided in ChOiRe? Our findings show that, firstly, with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions. Secondly, increasing the number of opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in section 3. Lastly, while including more opinions could possibly harm the performance of models, our answer consistency strategy enables LLMs to achieve the best results across three different K values.
Model ChatGPT ChatGPT-Inst GPT-4 Mistral
% of ITA of DIO-LLMtop8 + FEA + CoO 0.61 1.32 9.71 0.00
DIO-LLMtop8 + FEA + CoO 51.90 53.01 59.02 54.21
% of ITA of DIO-LLMtop10 + FEA + CoO 0.12 1.01 5.44 0.00
DIO-LLMtop10 + FEA + CoO 51.55 52.74 58.88 53.88
% of ITA of DIO-LLMtop12 + FEA + CoO 0.00 0.66 3.12 0.00
DIO-LLMtop12 + FEA + CoO 51.60 52.31 59.11 52.96
ChOiRe 52.21 53.26 59.30 54.43
Table 4: Extra analysis on ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. ITA stands for "Impossible To Answer".
6.2 Error Analysis
FEA Misses Key Explicit Personae.
Despite showing promising results in removing unuseful explicit personae depicted in table 3, we observe that LLMs sometimes misselect relevant personae. One such example is the top-left of section E.6. We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
LLMtop-$K$ Opinions Include Less Relevant Ones.
While LLMs generally demonstrate a commendable ability to rank implicit opinions by usefulness, as exemplified in section E.2, we also observe they frequently include less relevant, or even irrelevant opinions to the ranked list such as in section E.6-bottom. We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
LLMs May Not Follow the Instructions.
Although ChatGPT and ChatGPT-Instruct demonstrate a robust ability to adhere to our instructions for opinion prediction via CoO, the same level of proficiency is not observed in Mistral and GPT-4, as shown in section E.6-top-right. We posit this disparity arises from the fact that ChatGPT and ChatGPT-Instruct excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.
7 Conclusion
We propose ChOiRe, a four-step solution framework for individual opinion prediction via differentiating the utilization of user’s explicit versus implicit personae. We further introduce Chain-of-opinion reasoning and answer consistency over variable numbers of input implicit personae guiding the models to derive thorough predictions. ChOiRe achieves strong SOTA results with limited inference calls, demonstrating its strong effectiveness. Additionally, Steps (i) and (ii) of ChOiRe significantly improve the fine-tuning of opinion-aligned models. We strongly suggest that our method should only be used for positive moral intents, avoiding making LLMs echo chambers Vicario et al. (2016). In the future, we will focus on developing frameworks that utilize explicit and implicit personae more efficiently.
Limitations
One limitation of our proposed ChOiRe framework is that it requires the LLMs to have a good capability in following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions one by one via CoO. However, we foresee that this limitation is going to be overcome by cutting-edge AI language models, in the present and near future. Additionally, our method also utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world. However, to what extent is the personal information provided, our ChOiRe is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions.
Ethical Considerations
Characterizing and predicting human opinions with LLMs can be directly applied to personalize and align machines to users’ values, and cultural beliefs. Nonetheless, there exist unwanted situations when LLMs with our techniques can be misused for unethical purposes and biased opinions.
Bias Amplification and Fairness.
A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Vicario et al. (2016). This can ultimately harm users by reinforcing polarized or undesirable views. To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed ChOiRe involves presenting user demography or ideology group responses alongside personalized answers. Additionally, CoO can encourage users to reflect on their previous viewpoints.
Privacy and Consent.
Users may not always be aware of or have control over the extent of personalization applied to the content they receive. Therefore, empowering users to have control over AI-generated opinions is essential. Users should be able to customize and adjust the explicit and implicit personae used for opinion prediction. This customization can help mitigate potential biases and provide individuals with AI-generated opinions that align more closely with their values and preferences.
Human Evaluation.
Through human evaluations, we observe that our proposed method does not generate any discriminatory, insulting responses. We validate the intermediate steps of our proposed ChOiRe by human evaluation which involves manual labor. We hire annotators to score, and the hourly pay is set to \$$\$15$, which is higher than the local statutory minimum wage. Therefore, we do not anticipate any major ethical concerns raising from human evaluations.
Appendix A Baselines
A.1 Baselines Implementation Details
Prompting.
[…]
Model ChatGPT/ChatGPT-Inst/GPT-4/Mistral
Guns Auto. & driverless vehicles Views on gender Com. types & sex. harassment Race
W/o persona 53.07/37.30/——-/30.48 47.73/48.26/——-/41.72 50.53/42.94/——-/37.39 47.73/41.67/——-/29.34 41.95/45.28/——-/37.55
DIO-top8 53.87/57.00/60.39/44.73 45.33/44.78/53.22/41.72 53.21/52.15/63.73/40.09 43.47/45.24/42.86/35.45 43.06/44.65/55.17/41.11
DIO-top8 + CoT 54.55/52.33/——-/55.48 47.22/46.77/——-/49.00 48.11/57.67/——-/54.28 42.39/42.26/——-/42.01 45.63/43.40/——-/49.78
DIO-top8 + SC 54.40/52.85/——-/56.57 43.73/48.26/——-/52.31 55.61/56.44/——-/56.30 45.33/40.48/——-/42.01 45.00/43.40/——-/50.00
ChOiRe 57.06/58.21/63.37/58.00 49.25/51.92/50.00/53.75 59.23/53.07/71.43/57.78 39.88/44.14/47.96/42.08 42.77/47.28/50.57/51.44
Gender & Leadership America in 2050 Trust in science Biomedical & food issues Misinformation
W/o persona 53.13/50.83/——-/43.51 39.73/39.13/——-/41.95 50.40/47.29/——-/48.34 53.87/53.63/——-/53.21 46.93/40.38/——-/53.63
DIO-top8 48.27/54.70/65.55/50.23 46.93/46.20/43.70/35.14 54.93/61.58/61.54/51.65 52.27/55.86/58.03/52.78 49.33/52.11/52.71/50.77
DIO-top8 + CoT 48.58/50.83/——-/55.79 43.05/48.91/——-/43.76 54.10/65.02/——-/58.28 56.91/57.54/——-/57.08 49.57/53.99/——-/53.19
DIO-top8 + SC 49.07/53.60/——-/57.87 45.87/47.83/——-/46.03 56.27/65.52/——-/58.94 53.07/57.54/——-/58.58 45.00/53.52/——-/53.85
ChOiRe 52.22/57.78/63.03/57.87 49.46/48.99/45.37/47.50 56.43/55.50/68.46/60.37 54.75/57.26/61.61/58.58 46.45/53.62/57.36/53.85
Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views
W/o persona 43.24/40.28/——-/33.64 47.06/44.36/——-/46.08 43.67/49.15/——-/34.07 46.13/46.71/——-/40.42 40.80/48.95/——-/46.20
DIO-top8 53.24/47.22/47.73/43.31 57.22/57.89/62.50/47.42 45.60/51.98/63.81/41.87 49.60/57.23/66.67/41.27 56.80/46.85/62.07/44.13
DIO-top8 + CoT 53.38/47.22/——-/56.91 59.57/55.64/——-/54.36 47.65/51.98/——-/51.45 46.42/56.58/——-/51.06 53.30/45.45/——-/50.80
DIO-top8 + SC 54.05/47.22/——-/58.06 55.35/54.89/——-/57.04 46.13/51.98/——-/52.76 46.42/55.26/——-/52.89 57.33/47.55/——-/51.67
ChOiRe 54.29/53.33/52.27/58.06 60.00/58.77/63.89/58.50 52.33/50.13/64.76/51.89 44.74/55.26/64.58/52.76 51.05/53.74/67.82/53.34
Table 5: Fine-grained accuracy results of ChatGPT/ChatGPT-Instruct/GPT-4/Mistral. DIO stands for Demographic + Ideology + Opinions (section 4).
Fine-tuning.
[…]
Appendix B Extra Analysis
B.1 Additional Baseline Comparisons
Model ChatGPT Mistral-7B-Instruct-v0.2
DIO-top8 50.22 44.16
DIO-top8 + FEA 50.64 44.99
DIO-top8 + Random FEA (S=2000) 49.47 42.23
DIO-top8 + Random FEA (S=2024) 48.85 43.36
DIO-LLMtop8 51.03 45.86
DIO + Random LLMtop8 (S=2000) 48.13 44.58
DIO + Random LLMtop8 (S=2024) 49.21 43.84
Table 6: Accuracy results of ChatGPT and Mistral with two trivial variants with two different random seeds 2000 and 2024 in section B.1.
In this section, we compare ChOiRe’s FEA and LLMtop-$K$ steps with two simple variants outlined in table 6. Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs? The first variant, DIO-top8 + Random FEA, involves randomly selecting $5/12$ and $9/12$ explicit persona attributes. The second variant entails randomly selecting $8$ implicit persona opinions instead of using ChOiRe’s LLMtop-$K$ step. From table 6, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes. Additionally, randomly selecting $8$ implicit persona opinions also adversely affects the models, particularly ChatGPT. These observations underscore the effectiveness and importance of ChOiRe’s FEA and LLMtop-$K$ steps.
B.2 Fine-grained Results of ChatGPT, ChatGPT-Instruct, and GPT-4
Table 5 presents the fine-grained results of ChOiRe and baselines for ChatGPT, ChatGPT-Instruct, and GPT-4.
B.3 ChOiRe versus Self-Consistency
Fig. 3 presents the improvements per topic comparision between ChOiRe and the SOTA baseline with Self-Consistency Wang et al. (2023b) DIO-top8 + SC. We observe that ChOiRe improves $11/15$ topics for both ChatGPT and ChatGPT-Instruct.
Figure 3: % of improvements over the SOTA method (DIO-top8 + SC) with ChatGPT-Instruct (left) and ChatGPT (right).
B.4 Accuracy Distribution over Users
Fig. 4 shows the accuracy distribution over users of ChOiRe with ChatGPT. We see that the peak accuracy is at 0.5 for the majority, with a few users scoring zero and over 20 achieving perfection.
Figure 4: Frequency distribution of accuracy over users by ChOiRe.
B.5 Top-3 Removed Explicit Personae Attributes
Table 7 presents the top-3 explicit personae that got removed the most by the LLMs. Among the removed personae, "Citizenship" appears to be the highest-frequency one across models, followed by "Race".
Topic ChatGPT ChatGPT-Instruct Mistral-7B-Instruct-v0.2
Guns ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Frequency of religious attendance’, ’Religion’ ’Citizenship’, ’Education’, ’Religion’
Automation & driverless vehicles ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Religion’, ’Frequency of religious attendance’
Views on gender ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Religion’, ’Frequency of religious attendance’
Community types & sexual harassment ’Citizenship’, ’Race’, ’Gender’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Education’, ’Race’, ’Political Party’
Biomedical & food issues ’Citizenship’, ’Race’, ’Marital status ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Marital status’
Gender & Leadership ’Citizenship’, ’Race’, ’Region’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Region’, ’Race’, ’Citizenship’
America in 2050 ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’
Trust in science ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Region’
Race ’Citizenship’, ’Marital status’, ’Age’ ’Citizenship’, ’Age’, ’Religion’ ’Marital status’, ’Education’, ’Age’
Misinformation ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Race’, ’Religion’
Privacy & Surveillance ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Religion’, ’Race’, ’Region’
Family & Relationships ’Citizenship’, ’Race’, ’Region’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Race’, ’Religion’
Economic inequality ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Gender’, ’Citizenship’, ’Religion’
Global attitudes ’Marital status’, ’Race’, ’Citizenship’ ’Citizenship’, ’Marital status’, ’Race’ ’Gender’, ’Frequency of religious attendance’, ’Marital status’
Political views ’Citizenship’, ’Marital status’, ’Frequency of religious attendance’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Frequency of religious attendance’, ’Gender’, ’Citizenship’
Table 7: Top-3 explicit personae that got removed the most by the LLMs.
B.6 Ranking Consistency for LLMtop-$K$ Step
We record the average Overlap coefficient Vijaymeena and Kavitha (2016) among $5$ ranking outputs from $5$ input strategies in fig. 5. The performance of those input strategies is further presented in table 8 on $300$ random samples.
Figure 5: ChatGPT and Mistral-7B-Instruct-v.02 overlap coefficient values for different values of $K$. We observe that for $K$ is large enough ($K\geq 8$), the coefficient value is relatively acceptable ($\geq 0.6$).
Model Method Semantic Similarity Order Seed = 2024 Seed = 5 Seed = 2000 Seed = 15 Std
ChatGPT DIO-LLMtop8 - 51.03 50.95 51.11 - 0.0652
Mistral-7B-Instruct-v0.2 DIO-LLMtop8 - 45.86 45.55 45.36 - 0.2060
Table 8: Accuracy results of ChatGPT and Mistral on our test set with DIO-LLMtop8 where different orders of input implicit persona opinions are tested for LLMtop-K step.
B.7 Kendall’s Tau Scores for Ranking Agreements
Figure 6: Left: Ranking agreements between ChatGPT top-$K$ and semantic similarity top-$K$. Right: Between ChatGPT-Instruct top-$K$ and semantic similarity top-$K$. One example that has a high disagreement score is shown in section E.2.
Figure 7: Ranking agreements between Mistral top-$K$ and semantic similarity top-$K$.
Fig. 6 shows our ranking agreement scores between ChatGPT and Semantic similarity metric (Left), and ChatGPT-Instruct and Semantic similarity metric (Right). We observe that the two ranking orders have minimal monotonous relations with means approximating 0 and low standard deviations. More specifically, with ChatGPT, the maximum agreement is 0.6000 while the minimum is -0.5895 and the Kurtosis is -0.2173. For ChatGPT-Instruct, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017.
B.8 Consistency Scores
Table 9 presents the exact consistency scores for the fig. 2. Besides CoO consistently outperforming CoT, we also observe that when the temperature is increased, the consistency score is decreased, which is intuitive.
Model Temperature Consistency Score (%)
DIO-top8 + CoT 0.3 84
DIO-top8 + CoO 0.3 86
DIO-top8 + CoT 0.6 79
DIO-top8 + CoO 0.6 82
DIO-top8 + CoT 0.9 58
DIO-top8 + CoO 0.9 60
Table 9: Consistency scores of CoT and CoO on $100$ random question-answer pairs. We sample $5$ answers per question and measure the % of questions that have all $5$ identical answers.
Appendix C Prompts and Prompts Analysis
C.1 Prompt Templates for Filtering Explicit Personae
We present the prompt template for selecting relevant explicit personae for answering the question below. The template is hand-crafted and we use Chain-of-Thought (CoT) prompting Kojima et al. (2022) via adding "answer the above question step by step".
A person can be described by the following attributes: {original_attribute_list} Based on the above list of demographic information above, now I give you a new question with possible answer choices: Question: ’{test_question}’ Answer choices: ’{test_choices}’ Please analyze which attributes in the demographic information are useful for you to answer the above question step by step. Give me the output in the Python list format: [...] Give me the answer in the format below: Explanations: ... Answer: [...]
C.2 Prompt Templates for Implicit Feature Ranking
We provide our hand-crafted prompt template for ranking implicit personae opinions in the usefulness order below:
Given social behavior question-answer pairs answered by a user about his/her opinions about {subtopic}: {original_persona_question_order} You are an expert in analyzing the social behaviors of a user. Given a new question asking him/her: ’{test_question}’ Your task is to sort the list of given question-answer pairs in descending order such that the first question-answer pair brings the most useful information to answer the new question, whilst the last question-answer pair brings the least useful information. Give me the answer in the form of a Python list of indexes: Answer: [...]
C.3 Prompt Templates for Baselines Techniques
We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
C.3.1 W/o Persona Santurkar et al. (2023)
The W/o Persona prompt is provided below.
Question: {question} Answer choices: {choice} Complete the answer by the following format without any explanation: Answer: A. or B. or C. or D. or E...
C.3.2 DIO-top8 Hwang et al. (2023)
The DIO-top8 prompt is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, which answer choice is the user most likely to choose? Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E....
C.3.3 Self-refine Madaan et al. (2023)
The Self-refine prompts Madaan et al. (2023) are provided below, feedback step and refine step respectively.
You are given a question and an answer for that question. Analyze the question and the answer and provide some feedback on the answer to the question. Don’t change the answer, just provide feedback. Question: {test_question} Choices: {choices} Answer: {selected_choice} Feedback:
You are given a question, an answer to that question and a feedback to the answer. Based on the feedback, refine your answer and generate the final answer in around 170 words. Question: {test_question} Answer: {selected_choice} Feedback: {feedback} Refined answer: new_choice + explanation
C.3.4 Chain-of-Thought Kojima et al. (2022)
The CoT prompt template is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E.... Explanations:...
C.3.5 Chain-of-Opinion (Ours)
Our CoO prompt template is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step by explaining and analyzing each of the Opinions and Demographic Information: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E.... Explanations:...
C.4 Prompting Costs for API Models
Our prompting costs for API models are reported in table 10. We observe that for ChatGPT and ChatGPT-Instruct, ChOiRe costs around $7$ and $10$ more US\$ dollars in total compared to baseline DIO-top8 + SC. However, these extra amounts of costs are worth it because we gain significant improvements over all the baselines and especially huge improvements for some topics. Additionally, for GPT-4, ChOiRe costs a similar price with the baseline DIO-top8 while DIO-top8 + SC costs nearly double our price. This is because we perform the FEA and LLMtop-$K$ steps of ChOiRe by ChatGPT, which are relatively cheap.
DIO-top8 DIO-top8 + CoT DIO-top8 + SC ChOiRe Model
Ave. consumed #tokens 562.72 623.62 995.89 3142.86 ChatGPT
Total US\$ 3.01 3.73 6.82 13.95 ChatGPT
Ave. consumed #tokens 562.72 630.58 1019.31 3121.72 ChatGPT-Instruct
Total US\$ 3.12 3.84 7.11 19.99 ChatGPT-Instruct
Ave. consumed #tokens 559.27 - 1021.14* 3180.82 GPT-4
Total US\$ 91.19 - 226.15* 123.30 GPT-4
Table 10: Prompting cost analysis of ChOiRe and other baselines as of 1st Feb 2024. * denotes our estimation on $50$ random samples.
Appendix D Human Evaluation
D.1 Human Rating Instructions
Our details of human rating instructions are provided in table 11 for all the criteria. It is worth noting that selecting all features can’t get a high FEA Satisfaction score, according to our instructions. In addition, if the selected explicit personae fall among several scores, the annotators are instructed to take the minimum score.
Criterion Scoring Instruction
1: The number of filtered-out explicit personae that are directly relevant for answering the question is more than 3.
1: The number of selected explicit personae that are somewhat irrelevant for answering the question is more than 3.
2: The number of filtered-out explicit personae that are directly relevant for answering the question is 2 or 3.
FEA Satisfaction 2: The number of selected explicit personae that are somewhat irrelevant for answering the question is 2 or 3.
3: The number of filtered-out explicit personae that are directly relevant for answering the question is less than or equal to 1.
3: The number of selected explicit personae that are somewhat irrelevant for answering the question is less than 2.
1: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question is more than 4.
LLMtop-$K$ Satisfaction 2: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question from 2 to 4.
3: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question is less than or equal to 1.
1: The CoO has limited or flawed reasoning thoughts with inadequate support.
CoO Reasonableness 2: The CoO has some reasoning thoughts with decent support but room for improvement.
3: The CoO has strong, clear, and well-supported reasoning thoughts with a comprehensive understanding.
1: The generated CoO explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
CoO Follow the Instruction 2: The generated CoO explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
3: The generated CoO explanation follows perfectly the instruction via explaining all the explicit and implicit attributes one by one.
Table 11: Human rating instructions. FEA, LLMtop-$K$, and CoO stand for Filtering Explicit Personae Attributes, Implicit Personae Opinions Ranking, and Chain-of-Opinion reasoning (section 3).
Appendix E Examples
E.1 FEA Example with ChatGPT
Fig. 8 shows an FEA example with ChatGPT. We observe that by removing unnecessary explicit personae including "Age", "Citizenship", "Education", "Income", "Marital Status", "Race", "Frequency of religious attendance", ChatGPT predicts the opinion accurately, while without removing, an incorrect prediction was made.
Figure 8: FEA example with ChatGPT.
E.2 Example of High Disagreement between Rankings
Fig. 9 illustrates one example of the high disagreement between orders by semantic similarity scores and LLM (ChatGPT). We derive three observations, as discussed in section 6.1. First, not all top-8 opinions by semantic similarity scores help predict the opinion. For example, $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, while an incorrect prediction is made with the semantic similarity order.
Figure 9: Example of the high disagreement between orders by semantic similarity scores and LLM (ChatGPT).
E.3 Example of Inconsistent Answers Generated by CoT
Fig. 10 illustrates an example of the inconsistent answers generated by ChatGPT with Chain-of-Thought Kojima et al. (2022) (CoT). It is observed that different subsets of top-8 implicit personae opinions are mentioned in the two explanations, leading to varied final answers.
Figure 10: Example of the inconsistent answers generated by ChatGPT with Chain-of-Thought.
E.4 Example of Chain of Opinion Reasoning
Fig. 11 presents an example of the answer generated by ChatGPT using Chain of Opinion (ours) versus Chain of Thought Wei et al. (2022) prompting methods.
Figure 11: Example of an answer generated by Chain of Opinion versus Chain of Thought prompting with ChatGPT.
E.5 Example of Answer Consistency with Dynamic Numbers of Opinions
Fig. 12 shows an example of the answer generated by GPT-4 using Chain of Opinion (ours) reasoning with different numbers of provided historical opinions.
Figure 12: Example of our answer consistency technique (ours), generated by GPT-4.
E.6 Error Analysis Examples
Fig. 13 illustrates our error analysis examples of ChOiRe with ChatGPT. The top-left frame is an example of FEA missing key explicit personae. The bottom one is an instance demonstrating the error of the LLMtop-$K$ algorithm including less relevant opinions. The top-right rectangular is an example from GPT-4, showing that it does not follow human instructions to predict opinion via chain-of-opinion reasoning.
Figure 13: Error analysis examples of ChOiRe with ChatGPT.

### v5

@@
-ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning
+Aligning Large Language Models with Human Opinions through Persona Selection and Value–Belief–Norm Reasoning
 Abstract
-Warning: This paper includes examples that may be deemed sensitive or offensive.
-Aligning language models (LMs) with human opinion is challenging yet vital to enhance their grasp of human values, preferences, and beliefs.
-We present ChOiRe [^1] Codes and data will be available at https://github.com/dxlong2000/ChOiRe., a four-step framework to predict human opinion which differentially models the user’s explicit personae (i.e.
-demographic or ideological attributes) that are manually declared, and implicit personae inferred from user historical opinions.
-ChOiRe consists of (i) an LM analyzing the user’s explicit personae to filter out irrelevant attributes; (ii) the LM ranking the implicit persona opinions into a preferential list; (iii) Chain-of-Opinion (CoO) reasoning, where the LM sequentially analyzes the explicit personae and the most relevant implicit personae to perform opinion prediction; (iv) and where ChOiRe executes Step (iii)’s CoO multiple times with increasingly larger lists of implicit personae to overcome insufficient personae information to infer a final result.
-ChOiRe achieves new state-of-the-art effectiveness with limited inference calls, improving previous techniques significantly by 3.22%.
-We also show that ChOiRe’s Steps (i) and (ii) can significantly better fine-tune opinion-aligned models, by up to 18.44%.
+Reasoning and predicting human opinions with large language models (LLMs) is essential yet challenging.
+Current methods employ role-playing with personae but face two major issues: LLMs are sensitive to even a single irrelevant persona, changing up to 30% of the predictions; and LLMs fail to reason strategically over personae.
+We propose Chain-of-Opinion (COO [^1] Our codes and data will be made publicly available here.), a simple four-step solution modeling which and how to reason with personae, inspired by the Value–Belief–Norm (VBN) theory.
+COO differentiates between explicit personae (demographics and ideology) and implicit personae (historical opinions), involves: (1) filtering irrelevant attributes from explicit personae; (2) ranking implicit personae into a preferential list for selecting top-k; (3) applying novel VBN reasoning to extract user environmental and personal value, belief, and norm variables for accurate and reliable predictions; and (4) iterating VBN reasoning with progressively larger lists of implicit personae to handle potential persona insufficiency.
+COO efficiently achieves new state-of-the-art opinion prediction via prompting with only 5 inference calls, improving prior techniques by up to 4%.
+Notably, fine-tuning LMs with COO’s data results in significantly better opinion-aligned models, by up to 23%.
+Aligning Large Language Models with Human Opinions through
+Persona Selection and Value–Belief–Norm Reasoning
+Do Xuan Long1,2, Kenji Kawaguchi1, Min-Yen Kan1, Nancy F.
+Chen2 1National University of Singapore, 2Institute for Infocomm Research (I2R), A*STAR xuanlong.do@u.nus.edu, {kenji,knmnyn}@nus.edu.sg, nfychen@i2r.a-star.edu.sg
 1 Introduction
-Language models (LMs) are becoming indispensable tools, serving in various roles such as dialogue agents OpenAI (2022); Google (2022), data analysts Wang et al.
-(2023a); Cheng et al.
-(2023), and decision support Ye et al.
+Pre-trained large language models (LLMs) are becoming indispensable tools, serving in various assistant roles such as dialogue agents OpenAI (2022); Google (2022) and data analysts Cheng et al.
 (2023).
-LMs also demonstrate the capability to model distinct opinions which influence response generation on input queries Bai et al.
+Notably, they demonstrate the capability to model distinct opinions that influence response generation Bai et al.
 (2022); Glaese et al.
@@
 (2023).
-Unfortunately, the opinions modeled by language models are shaped by the extensive training and feedback data, which are themselves influenced by countless human perspectives, making them inherently challenging to model.
-As human–AI interactions become common, it becomes imperative to align models with human opinion to meet individual and personalized expectations.
-Alignment frameworks such as RLHF Christiano et al.
+Unfortunately, their opinions are shaped by extensive training data, which are themselves influenced by countless human perspectives and thus challenging to comprehend.
+As human–AI interactions grow, instructing models to reason in alignment with human opinions is crucial for effective personalization.
+Although fine-tuning alignment methods such as RLHF Christiano et al.
 (2017); Ouyang et al.
-(2022) have been devised to train and fine-tune LMs.
-Nonetheless, applying them to align large language models (LLMs) with human opinions is difficult because they require both significant computational resources and high-quality supervised feedback data, which is itself challenging to collect.
-Therefore, prompt-based opinion alignment through persona has been studied as an easy-to-deploy and easy-to-combine resource-parsimonious alternative Perez et al.
-(2023); Simmons (2023); Santurkar et al.
-(2023); Deshpande et al.
+(2022) are widely employed, their application to personalized opinions remains challenging due to the significant compute and data required.
+Prompt-based role-playing frameworks using personae have emerged as alternatives.
+Early work in this direction focused on aligning models with social groups rather than individuals Perez et al.
+(2023); Santurkar et al.
 (2023).
-However, even when aligning with well-represented groups, persona-based methods result in low steerability Santurkar et al.
+However, Santurkar et al.
+(2023) found that simple persona-based prompting exhibits low steerability, even for well-represented groups.
+Further, Hwang et al.
+(2023) revealed significant opinion variation among individuals with similar demographics, underscoring the challenge of aligning LLMs to individual opinions.
+They subsequently proposed a naïve solution to model individuals by incorporating the user’s demographics and ideology (we term these explicit personae), alongside their historical opinions (implicit personae).
+While naïvely incorporating explicit and implicit personae into the prompt shows promise for individualization Santurkar et al.
+(2023); Hwang et al.
+(2023), this approach is suboptimal.
+Both persona types include ones that may be irrelevant to the opinion of interest, such as the ‘‘Citizenship’’ for the ‘‘Gun’’ question in Fig.
+1.
+This poses a problem as we observe that LLMs are highly sensitive to such noise (detailed in Section 3), highlighting the task of relevant persona selection as important yet unsolved.
+Moreover, effective reasoning over explicit and implicit personae for opinion prediction is challenging: we find that Chain-of-Thought Wei et al.
+(2022); Kojima et al.
+(2022) unexpectedly fails to improve opinion prediction in LLMs, due to reasoning inconsistencies.
+In line with the Value-Belief-Norm (VBN) theory Stern et al.
+(1999) which asserts that values, beliefs, and norms influence human behavior distinctly, we argue that the explicit and implicit personae should be processed and utilized differently.
+The explicit offers clear insights into user environmental values (binary relevancy), while the implicit reveals nuanced and context-specific beliefs and norms shaped by personal experiences (complex relevancy).
+Building on this, we introduce Chain-of-Opinion (COO, Fig.
+1), a novel four-step framework that optionally leverages both persona types to address the above challenges: (1) an LLM analyzes explicit personae to discard irrelevant ones; (2) the LLM ranks implicit persona opinions by usefulness and selects the top-$K$; (3) VBN reasoning where the LLM generates high-level environmental values from selected explicit personae and nuanced individual beliefs and norms from top-K implicit personae before deriving a prediction; and (4) COO applies VBN reasoning using varying $K$ of implicit personae from step (2) to prevent the model from refusing to answer due to insufficient personae.
+COO achieves state-of-the-art opinion prediction with persona-based prompting by just five inference calls (Section C.1).
+Moreover, fine-tuning with data from COO’s steps (1–3) enhances LMs by up to 23%, resulting in Flan-T5 base model Chung et al.
+(2024) comparable with GPT-4 OpenAI (2023b).
+COO is highly generalizable in scenarios with missing personae (Section 6.1), and its four steps can motivate and be applicable to other personalized tasks involving explicit personae and user historical views (Section B.8).
+2 Related Work
+LLM role-play with personae.
+Aligning language models with human behavior via personae is a growing study area.
+Such alignment increases user satisfaction and personalization Wang et al.
+(2023c); Chen et al.
 (2024).
-This raises significant concerns when using LMs to model for individual users.
-Since individuals hold nuanced opinions that evolve and are influenced by situational factors, it is more challenging to achieve individual opinion LM alignment.
-Hwang et al.
-(2023) find significant opinion variations among individuals sharing the same demographics, exposing flaws in current group-focused alignment.
-They argue for individualized models, introducing an approach (detailed below) integrating user’s demographic and ideological attributes, (we term as explicit personae) and historical opinions (implicit personae) into the prompt for opinion prediction.
-While Hwang et al.
-(2023)’s naïve integration achieves promising results, we argue that it suffers from a few key limitations.
-First, it employs all explicit personae.
-However, we contend that only a subset is needed for accurate opinion prediction; including non-relevant personae may act as noise, harming predictive performance.
-Second, Hwang et al.
-(2023) utilize the top-$K$ semantically similar opinions with the question (here termed top-$K$ implicit personae).
-This approach is inefficient, as such similar opinions may not offer the most valuable information for opinion prediction.
-Additionally, our empirical experiments suggest that LMs may lack sufficient personae evidence with a fixed $K$ — dynamically adjusting $K$ per task can overcome such deficiencies.
-Finally, while Chain-of-Thought (CoT; Wei et al.
-2022; Kojima et al.
-2022) enables LMs to perform multi-step reasoning tasks effectively, we find that the naïve application of CoT does not help modern LLMs like ChatGPT with opinion alignment [^2] We validate all our claims on these limitations in Sections E.1, 6.1, 4 and 5, respectively..
-To address the above challenges, we propose ChOiRe [^3] Chain of Opinion Reasoning, pronounced as the English word “choir”.(fig.
-1), a four-step solution for opinion prediction leveraging LLMs’ strong data evaluation and analytic capabilities Wang et al.
-(2023a); Cheng et al.
-(2023).
-First, an LLM analyzes a target user’s explicit personae to discard irrelevant ones.
-Second, the LLM ranks implicit persona opinions in order of usefulness, selecting the top-$K$ as the most valuable.
-This surpasses the constraint of using semantic similarity scores.
-Third, we introduce Chain-of-Opinion (CoO), a designed variant of CoT that allows the LLM to explain and analyze selected explicit personae and top-$K$ implicit personae sequentially.
-ChOiRe applies self-consistency over CoO to provision the appropriate amount of user information for opinion inference.
-ChOiRe achieves new state-of-the-art (SOTA) in opinion alignment effectiveness and reliability, while using a limited inference budget.
-We conduct a thorough analysis to verify our hypotheses concerning explicit and implicit personae and defend our Chain-of-Opinion reasoning methodology.
-Our analysis further reveals that ChOiRe’s first two steps in handling explicit and implicit personae also helps to fine-tune opinion-aligned models.
-2 Related Work
-Aligning LMs with Humans.
-Aligning language models with human behaviour is a recent area of study as alignment can increase user experience satisfaction and utility Wang et al.
-(2023c).
-One line of work develops prompting techniques with user demographic information (e.g., political identity) to encourage LMs to output human-like responses.
+One line of work develops prompting techniques with user demographics, encouraging LLMs to output human-like responses.
 Argyle et al.
-(2023) show that by properly conditioning LMs with targeted identity and personality profiles, it is possible to produce biased outputs that strongly correlate with human responses.
-Furthermore, Simmons (2023) claims that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases.
-Despite recent advances, Santurkar et al.
-(2023) discovered that LMs align poorly with human opinions, as evidenced by model performance on public opinion polls.
-Hwang et al.
-(2023) recently propose to incorporate explicit and implicit personae to predict human opinions in new contexts.
-In section 1, we argue that this naïve strategy is suboptimal.
-ChOiRe overcomes these limitations.
+(2023) showed that by properly conditioning LLMs with targeted identity profiles, they produce biased outputs that strongly correlate with human responses.
+Furthermore, Simmons (2023) claimed that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases.
+Nevertheless, Santurkar et al.
+(2023); Hwang et al.
+(2023) discovered that LLMs align poorly with human opinions, as evidenced by model performance using explicit and implicit personae on public opinion polls.
+We argue that this strategy is suboptimal (Section 1) due to noisy personae and the inefficient reasoning strategy employed.
+COO overcomes these limitations.
 Reasoning with LMs via prompting.
@@
 (2023); Touvron et al.
-(2023) have enabled large language models (LLMs) to excel at various NLP tasks using zero- or few-shot prompting Liu et al.
+(2023) have enabled LLMs to excel at various reasoning NLP tasks via prompting Wei et al.
+(2022); Khot et al.
+(2023); Zhou et al.
+(2023); Wang et al.
+(2023b); Shinn et al.
 (2023).
@@
 (2022); Kojima et al.
-(2022) propose prominent Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps to solve multi-step reasoning tasks with higher fidelity and efficiency.
+(2022) proposed the popular Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps, aiding the solving of multi-step reasoning tasks with higher fidelity and efficiency.
 Can CoT analyze and predict human opinion effectively?
-We find that a naïve application of CoT does not help GPT-X models (section 5), but that an appropriate modification does.
-We propose Chain-of-Opinion (CoO) reasoning (section 3) that overcomes CoT’s limitations in this task.
-Noting that prompting techniques such as task decomposition Khot et al.
-(2023); Zhou et al.
-(2023) and retrieved-based methods Yao et al.
-(2023); Shinn et al.
-(2023) have been recently introduced, we focus only on the reasoning explanation aspect here given the abstractive and challenging nature of the task.
-Figure 1: ChOiRe overview, consisting of the four main steps (cyan background), as detailed in section 3.
-3 ChOiRe: A Chain of Opinion Framework
-Task Formalisation.
-We follow Santurkar et al.
-(2023), and formulate the opinion prediction task as multiple-choice question answering.
-Formally, a benchmark with $N$ data points is denotated as $D=\{\langle T,E,I,q,a\rangle_{n}\}_{n=1}^{N}$, where $T$, $E$ and $I$ indicate the (T)opic of a question $q$, the (E)xplicit personae and (I)mplicit personae of the user answering $q$ with opinion $a$.
-Following the prior work, $E$ consists of $12$ user demographic and ideology metadata attributes, and $I$ contains a number of the user’s historical opinions in the format of question–answer pairs.
-Models then learn to analyze $T,E,I,q$ and predict the opinion $a$.
+Surprisingly, a naive application of CoT fails to improve GPT-X models (Section 5.2).
+We attribute this to the reasoning inconsistencies of CoT and the complexity of the task: strategic reasoning is essential to consistently fully utilize the nuanced explicit and implicit personae.
+Our VBN reasoning (Section 4) overcomes CoT’s limitations.
+3 LLMs are Distracted by Irrelevant Personae
+Type Relevant +1 less/irr +3 less/irr +all less/irr
+persona relevant relevant relevant
+ChatGPT Explicit 37.56 35.53$\downarrow$ 34.51$\downarrow$ 34.35$\downarrow$
+Implicit 34.35 33.84$\downarrow$ 30.76$\downarrow$ 31.28$\downarrow$
+LLaMa 3.1 Explicit 26.39 24.36$\downarrow$ 23.85$\downarrow$ 22.84$\downarrow$
+Implicit 26.66 25.12$\downarrow$ 21.02$\downarrow$ 23.08$\downarrow$
+Table 1: The performances of ChatGPT (gpt-3.5-turbo-0125) and LLaMa3 (8B 3.1-it) significant drop (approx.
+1-6%) on Gun topic of OpinionQA Santurkar et al.
+(2023) when less or irrrelevant personae added.
+We study the sensitivity of LLMs to irrelevant personae which motivates COO in Section 4.
+This exploration is related to Shi et al.
+(2023) but we quantify the LLM sensitivity to personae instead.
+We find that LLMs can be easily distracted by explicit or implicit personae that are less or irrelevant to opinion questions.
+To examine this, we perform a semi-human evaluation to assess the relevancy of personae on 197 randomly chosen Guns questions from the OpinionQA dataset Santurkar et al.
+(2023).
+Each sample is denoted as $\{T,E,I,q,o,a\}$ where $T$, $E$, and $I$ indicate the topic, explicit personae (demographics and ideology), and implicit personae (historical opinions) of the user answering $q$ with opinion options $o$ and correct label $a$.
+To assess the relevance of each explicit persona in $E$ to $q$, we employ two native English undergraduates and ChatGPT (gpt-3.5-turbo-0125) OpenAI (2022).
+The annotators carefully examined ${q,a}$ and labeled 12 attributes of $E$ as relevant or irrelevant, determined by majority vote.
+The annotators achieve a good agreement of 60.2% Krippendorff’s alpha Krippendorff (2011).
+For implicit personae ($n=\sim 20$), assessing their relevance to $q$ by manual means is costly.
+Therefore, we compute the semantic similarity between them and $q$ using OpenAI text-embedding-ada-002 and label the top-8 as relevant and the rest as irrelevant.
+We test four different setups for each type of personae: (i) using only relevant ones; (ii) including one, (iii) three, or (iv) all irrelevant ones.
+To ensure that our results are consistent, we use Self-Consistency Wang et al.
+(2023b) with $5$ times sampling.
+We experiment with both representative closed and open source LLMs: ChatGPT and LLaMa3 Touvron et al.
+(2023).
+Results in Tab. 1.
+Surprisingly, adding a single irrelevant explicit persona results in a prediction change of $30\%$ for ChatGPT and $40\%$ for LLaMa, with $2\%$ performance drop for both.
+Meanwhile, adding one irrelevant implicit causes much smaller drops for both.
+Explaining this phenomenon, we have two hypotheses: first, both models rely more on explicit personae for opinion prediction; second, the so-called irrelevant implicit persona may still hold some relevance, as low semantic similarity does not entirely equate to low relevance (Appx.
 Fig.
-1 shows an overview of ChOiRe, consisting of four main steps (marked with a cyan background).
-First, ChOiRe employs an LLM to analyze and select a subset of relevant explicit personae, denoted as $E^{rel}\subseteq E$ for answering the opinion question $q$.
-The LLM then assesses the informativeness of the implicit personae ($I$) in predicting $q$, selecting the top-$K$ implicit personae (termed LLMtop-$K$).
-Next, an LLM is prompted to explain the provided explicit $E^{rel}$ and implicit LLMtop-$K$ personae sequentially in a Chain-of-Opinion (CoO) reasoning strategy.
-Finally, ChOiRe calls the LLM to predict the opinion $a$ with varying values of $K$ for the top-$K$ implicit personae.
-ChOiRe chooses the opinion with the highest frequency as the final prediction We detail the steps below.
-3.1 Filtering Explicit Personae Attributes (FEA)
-Accounting for explicit personae, which consist of the demographic and ideological metadata attributes of users — such as their age, income, and political ideology — has been shown to help models characterize and predict human opinions more accurately Hwang et al.
-(2023).
-However, which personae matter and which do not?
-are still open questions.
-Section E.1 shows such an example in full, where, when considering all of explicit personae, the model makes an incorrect prediction while removing unnecessary personae the model made a correct prediction.
-This may be caused by the LLM’s attention mechanism’s forcing the model to attend to all input tokens, even irrelevant ones.
-To filter out unnecessary explicit personae, we ask the LLM to reason and analyze how each persona is helpful for the model to predict the opinion via Chain-of-Thought to output a list of which personae are relevant given the question and the opinion answer choices [^4] We provide our FEA prompt in Section C.1.
-Surprisingly, we find that LLMs evaluate more than half of the explicit personae as irrelevant on average.
-We further conduct human evaluations to verify this finding in section 4.
-3.2 Implicit Personae Opinions Ranking (LLMtop-$K$)
-LLMs have been sensitive to selected demonstrations and their order in the prompts Perez et al.
-(2021); Luo et al.
-(2023); Gao et al.
-(2023).
-For predicting human opinions, we discover that LLMs are also sensitive to the chosen implicit personae opinions as input.
+6).
+Additionally, we observe that adding three/all irrelevant for both persona types significantly reduces performance by over 3-5% in absolute, indicating that irrelevant personae harm model predictions significantly.
+4 COO: A Chain of Opinion Framework
+These above findings suggest that for pre-trained LLMs, the choice of (relevant) personae as input is important and significantly impacts the model’s outcomes.
+Notably, we observe that training language models sees even greater benefits when incorporating only relevant personae.
+We introduce a simple and cost-efficient framework, termed Chain-of-Opinion (COO) characterizing explicit and implicit personae for opinion prediction.
+COO serves as an intermediate data preprocessing step for fine-tuning and prompting.
+We assume optional access to the user’s explicit and implicit personae.
+Let ${G}_{\mathcal{M}}:\mathcal{V}^{*}\to\mathcal{V}^{*}$ be the generation function of $\mathcal{M}$ where $\mathcal{V}$ is the model vocabulary, $Q$ be the concatenation of the test question and its answer choices, and $a$ be the correct label.
+4.1 COO in Prompting
+As introduced in Section 1 and depicted in Fig.
+1, COO distinctly processes explicit and implicit personae, motivated by the Value–Belief–Norm theory Stern et al.
+(1999), and consists of four main steps: Step 1.
+Filtering explicit personae; Step 2.
+Ranking implicit personae; Step 3.
+Value-Belief-Norm (VBN) reasoning; and Step 4.
+Answer consistency with dynamic numbers of opinions.
+Figure 1: COO overview with four main steps marked with the nuts. It processes explicit and implicit personae parallelly to facilitate the missing personae scenarios.
+Step 1. Filtering explicit personae (FEA).
+This step aims to binarily filter explicit personae (the set $E$, denoted in Section 3), including user demographics and ideology, for prediction.
+Irrelevant ones can harm the model performance significantly (Section 3), possibly because of LLM’s attention mechanism forcing the model to attend to all input tokens, including irrelevant ones.
+Since the relevancy of an explicit persona to a test question is apparent, we binarily filter out irrelevant explicit personae by instructing the LLM $\mathcal{M}$ to reason and analyze how each of them is helpful for the model to predict the opinion via Chain-of-Thought Wei et al.
+(2022):
+$E_{rel}:=G_{\mathcal{M}}([E,Q])$ (1)
+The instruction for $\mathcal{M}$ is skipped in Eq.
+1.
+Surprisingly, LLMs evaluate more than half of the explicit personae as irrelevant on average.
+We conduct human evaluations to examine this in section 5, and provide an example in full in Section E.2 showing when considering all explicit personae, the model yields an incorrect prediction while removing unnecessary personae, it offers a correct one.
+Step 2. Implicit personae opinions ranking (LLMtop-$K$).
+This step focuses on ranking and selecting implicit personae ($I$) consisting of user historical opinions for prediction.
+Identifying the most supportive ones for a test question is critical yet complex and often requires significant comprehension efforts.
 Hwang et al.
-(2023) rank the implicit personae opinions via semantic-similarity scores and select top-$K$.
-This strategy is suboptimal because the top-ranked opinions in terms of semantic similarity may not be the ones that provide the most supportive information for the models to predict opinions (section E.2).
+(2023) utilized semantic-similarity between ($q$, $i$) to select implicit persona $\forall i\in I$.
+This strategy is suboptimal because the top semantic similarity opinions may not be the ones that provide the most supportive information for the models (Section E.2, also we hypothesized in Section 3).
 As LLMs are shown to be good data analysts Wang et al.
 (2023a); Cheng et al.
-(2023), we propose to address the above challenge by utilizing LLMs to analyze and rank the implicit personae opinions in usefulness descending order.
-Our finding is that despite the output rankings from LLMs varying with different input orders of implicit personae opinions, the sets of LLMtop-$K$ opinions overlap by a good coefficient when $K$ is large enough ($\geq 8$) (fig.
+(2023), we propose to utilize LLMs to analyze and rank the implicit personae opinions in usefulness order:
+$I_{rel}:=top_{K}(G_{\mathcal{M}}([I,Q]))$ (2)
+The instruction of $\mathcal{M}$‘ is skipped in Eq.
+2.
+We fix $K$ when selecting top-$K$ implicit personae rather than having LLMs directly select useful ones like COO’s Step 1 because implicit personae involve more complex and nuanced information with varying degrees of relevance, unlike the binary relevance of explicit personae (Section 1).
+Additionally, we input $I$ to LLMs in a random order to enhance the versatility of our method.
+We validate this approach in Section 6.
+Our method prioritizes supportiveness in reasoning about opinions, unlike conventional demonstration selection focusing on semantic similarity or diversity Xu et al.
+(2024c).
+Step 3. Value-Belief-Norm reasoning (VBN).
+How can we best utilize $E_{rel}$ and/or $I_{rel}$ for opinion prediction?
+Popular prompting methods for reasoning, such as few-shot and zero-shot Chain-of-Thought (CoT) Wei et al.
+(2022); Kojima et al.
+(2022), typically instruct LLMs to reason step-by-step.
+However, for opinion prediction, a general step-by-step guide presents two key issues.
+First, the generated reasoning steps can be inconsistent, causing divergent outcomes, particularly at medium-high decoding temperatures ($\geq$ 0.6) (proven in Section 6.2 and Section E.2).
+This significantly undermines the reliability and accuracy of models.
+Second, step-by-step reasoning lacks interpretability, as it remains unclear how models process and prioritize multiple explicit and implicit personae.
+When numerous opinions and personae are input, it is difficult to discern which were used in forming the prediction and which were disregarded.
+[…]
+[…]
+$\rightarrow\hat{a}:=G_{\mathcal{M}}([E_{rel},I_{rel},Q])$ (4)
+Eq. 3 is achieved in a chain of thoughts:
+$(I_{1})$ Analyze the user’s demographics and ideology one by one to infer their social and environmental values.
+$(I_{2})$
+Analyze the user’s historical opinions one by one to infer their beliefs and norms from their social and environmental values.
+$(I_{3})$
+Which opinion is the user likely to choose?
+Fig.
+1 shows an example of VBN reasoning.
+Note that the instruction $(I_{1})$/$(I_{2})$ is skipped if $E_{rel}$/$I_{rel}$ is unavailable.
+Our human evaluation in Section 5.2 confirms LLMs are capable of reasoning about human values, beliefs, and norms.
+This reasoning offers two notable advantages.
+First, for each question, we ensure that the model explains and analyzes the provided personae one by one thoroughly without omitting any, resulting in more accurate predictions.
+Second, this method helps the model to output more consistent reasoning explanations, enhancing model reliability (Section 6.2).
+Step 4. Answer consistency with dynamic numbers of opinions.
+By fixing $K$ in the $LLMtop-K$ step, we observe that models such as GPT-4 OpenAI (2023b) may refuse to answer the question (Tab.
 5).
-Therefore, we propose to input the implicit personae opinions to LLMs in a random order to make our method more versatile.
-We also examine the case where we input the opinions in the semantic similarity order.
-We illustrate the prompt template in section C.2.
-By performing this step, our proposed method supports the usefulness of opinions in predicting the test opinions, rather than the semantic similarity.
-We term this method as LLMtop-$K$.
-3.3 Chain-of-Opinion Reasoning (CoO)
-Wei et al.
-(2022); Kojima et al.
-(2022) introduce few-shot and zero-shot Chain-of-Thought (CoT) prompting strategies demonstrating that by reasoning step-by-step, LLMs can achieve promising results on complex tasks.
-However, the sampled reasoning steps can be inconsistent, leading to possibly different outcomes Wang et al.
-(2023b).
-Furthermore, it is little known how the models perceive multiple implicit personae opinions, especially when many opinions are input, which one(s) the models used, which one(s) they didn’t for predicting the opinion?
-Our preliminary experiments with CoT (sections 6.1, E.4 and E.3) reveal that the CoT explanations can vary frequently based on different subsets of opinions mentioned in their explanations, leading to diverse final answers, especially when the decoding temperature is relatively high555greater or equal to 0.6.
-To mitigate this issue, we propose to instruct the LLMs to analyze the given explicit and implicit personae one by one before concluding the prediction via simply adding "explaining and analyzing how each of the Opinions and Demographic Information supports the question" into the prompt instruction.
-Given an LLM that can follow human instructions well such as ChatGPT OpenAI (2022), this addition offers two notable advantages.
-First, for each question, we ensure that the model explains and analyzes the provided personae one by one without missing any, possibly resulting in more thorough predictions.
-Second, this method helps the model to output more consistent reasoning explanations, enhancing its reliability (section 6.1).
-3.4 Answer Consistency with Dynamic Numbers of Opinions
-Prior work Hwang et al.
-(2023) fixes the number of implicit personae opinions for prediction to $K=8$.
-However, this approach occasionally results in models generating "...the answer cannot be determined." (tables 4 and E.5).
-We attribute this to insufficient user implicit personae opinions provided.
+We attribute this to insufficient implicit personae provided.
 Inspired by Self-Consistency (SC) Wang et al.
 (2023b), our approach involves sampling multiple answers using different $K$ values for a given question.
-The most frequent answer, along with its explanation, becomes the final prediction.
-Our method is distinct from SC since SC samples multiple answers with a fixed prompt.
-We experiment with $K\in\{8,10,12\}$ for efficiency.
-4 Evaluation
+The most frequent answer, along with the explanation of the first correct answer, becomes the final prediction.
+Our method is distinct from SC which samples multiple answers with a fixed prompt.
+We only use three values of K $\{8,10,12\}$ for efficiency.
+4.2 COO in Fine-tuning
+During fine-tuning, we adopt COO’s Steps 1, 2, 3 in Section 4.1 to create COO data.
+Each sample is denoted as {$E_{rel},I_{rel},\text{Env.
+Values},\text{Beliefs and Norms},Q,a$} where the model learns to predict $a$ given the rest variables.
+Since these steps require LLMs with strong instruction-following capabilities, we use ChatGPT as the COO data processor.
+5 Experiment
+5.1 Experimental Settings
 Dataset.
 We experiment on OpinionQA dataset Santurkar et al.
-(2023) — the only opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups.
-It covers $60$ US demographic groups, with $15$ topics, each comprising around $100$ questions, gathered from $5,340$ users.
+(2023) — the only open-sourced opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups.
+We also note the MRFP dataset Sun et al.
+(2023), which, due to privacy concerns, is unavailable for public access.
 Dataset preprocessing.
-Due to limited resources, we randomly sample $25$ users per topic for our experiments.
-For each user, we follow Hwang et al.
-(2023) to use $20\%$ of the implicit questions as the implicit persona.
+Due to limited resources, we follow Hwang et al.
+(2023) to sample a subset of OpinionQA for evaluation.
+We randomly select $25$ users per topic for our experiments.
+We use $20\%$ of the implicit questions for each user as the implicit persona.
 For the remaining $80\%$ implicit questions, we randomly select a maximum of $15$ implicit questions for testing.
 Our sampling method results in a total of $375$ users and $5,603$ implicit evaluation question–answer pairs.
-Our subset is highly representative because we gather 25 users from every topic and 15 questions per user.
-Rigorous statistical tests further validate the significance of our results which align closely with Hwang et al.
-(2023) testing on a larger subset using InstructGPT.
-Prompting Baselines.
-We use both closed-source ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b), and open-source Mistral-7B-Instruct-v.02 Jiang et al.
-(2023) as our LLMs, and compare ChOiRe with $5$ prompting methods: (1) W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; (2) Demographic + Ideology + top8 Opinions (termed DIO-top8), introduced by Hwang et al.
-(2023) demonstrating that integrating explicit and implicit personae enhances user opinion modeling and prediction, achieving state-of-the-art results on OpinionQA at that time; (3) DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al.
-(2022) version of DIO-top8 involving appending "answer the following question step-by-step" to prompts, aiming to explore whether CoT improves model performance in this task; (4) DIO-top8 + SC is the baseline which we apply the Self-Consistency technique with CoT Wang et al.
-(2023b) to DIO-top8 to select the most frequent answer generated by the model as the final opinion prediction; (5) DIO-top8 + Self-refine Madaan et al.
+Our subset is highly representative because we gather users from every topic and rigorous statistical tests further validate the significance of our results.
+Baselines.
+For prompting experiments, we use both closed- and open-source LLMs: ChatGPT OpenAI (2022), ChatGPT-it OpenAI (2023a), GPT-4 OpenAI (2023b), and Mistral-7B-it-v.02 Jiang et al.
+(2023) where each model performs all the COO’s steps.
+We compare COO with $5$ prompting methods: W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; Demographic + Ideology + top8 Opinions (DIO-top8), introduced by Hwang et al.
+(2023) which achieves state-of-the-art results on OpinionQA at that time; DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al.
+(2022) version of DIO-top8 by appending "answer the following question step-by-step" to prompts; DIO-top8 + SC is when we apply the Self-Consistency Wang et al.
+(2023b) with CoT to DIO-top8; DIO-top8 + Self-refine Madaan et al.
 (2023) interactively feedbacks and refines the answers by LLMs.
-We do not experiment with InstructGPT Ouyang et al.
-(2022) like Hwang et al.
-(2023) since this model is going to be deprecated and replaced by ChatGPT-Instruct.
-For GPT-4, we only run the main experiment and we use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget.
-All the prompts and costs are in appendix C, implementation details in section A.1, and more baselines in section B.1.
-Fine-tuning Baselines.
-We further investigate whether ChOiRe’s FEA and LLMtop-$K$ steps (section 3) also improve fine-tuning for opinion-aligned models.
-We first create the fine-tuning data by using ChatGPT to perform ChOiRe’s FEA and LLMtop-$K$ steps ($K=8$) on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones.
+For GPT-4, we only run the main experiment and use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget.
+We provide all the prompt templates and cost analysis in Appendix C, implementation details in Appendix A, and more baselines in Section B.2.
+For fine-tuning, we use ChatGPT to perform COO’s steps 1, 2 ($K=8$), 3 as noted in Section 4.2 on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones.
 We then fine-tune and evaluate GPT-2 models (base, large) Radford et al.
@@
 (2024).
-Fine-tuning details are provided in section A.1.
-Metrics.
-We employ Accuracy and Collapsed Accuracy as the automatic evaluation metrics following Hwang et al.
+The details are shown in Appendix A.
+Prompting Fine-tuning
+Model ChatGPT ChatGPT-it Mistral-7B-it-v0.2 GPT-2-base GPT-2-large FlanT5-base FlanT5-large
+W/o personae 46.60 / 65.72 44.91 / 63.60 41.24 / 59.54 36.28 / 52.62 21.94 / 39.11 48.98 / 68.33 39.83 / 58.43
+[…]
++ Self-refine 43.14 / 65.33 42.71 / 62.98 36.23 / 55.06
++ CoT 49.96 / 69.05 51.90 / 71.51 52.25 / 71.95
++ CoT-SC 50.58 / 69.66 52.06 / 71.87 53.14 / 72.88
+[…]
++ VBN (Step 3) 51.38 / 70.32 52.61 / 71.90 53.59 / 73.46 24.79 / 43.50 28.73 / 47.09 58.21 / 77.80 56.87 / 76.18
+DIO-LLMtop8 (Step 2) 51.03 / 70.31 52.80 / 72.60 45.86 / 64.98 22.65 / 41.12 28.86 / 47.60 57.97 / 77.46 58.20 / 77.56
++ FEA (Step 1) 51.19 / 70.69 52.97 / 72.84 45.23 / 64.73 25.05 / 44.41 29.54 / 48.66 57.45 / 77.13 59.00 / 78.46
++ FEA + VBN 52.16 / 71.90 53.08 / 72.92 54.56 / 74.37 26.17 / 45.92 30.21 / 49.63 59.62 / 78.87 60.13 / 78.92
+COO (ours) 52.66$\dagger$ / 72.75$\dagger$ 53.58$\dagger$ / 73.80$\dagger$ 54.40$\dagger$ / 74.26$\dagger$ 26.17$\dagger$ / 45.92$\dagger$ 30.21$\dagger$ / 49.63$\dagger$ 59.62$\dagger$ / 78.87$\dagger$ 60.13$\dagger$ / 78.92$\dagger$
+% w/ best baseline + 4.11 / + 4.43 + 2.91 / + 2.68 + 2.37 / + 2.57 + 23.26 / + 18.84 + 21.13 / + 17.55 + 8.40 / + 5.18 + 9.45 / + 5.52
+Table 2: Accuracy (Acc) / Collapsed Accuracy (CAcc) experimental results.
+FEA, LLMtop8, and VBN are COO’s Steps 1 (explicit), 2 (implicit), and 3 (reasoning).
+$\dagger$ denotes our method outperforms baselines with p-value < 0.01 under t-test (Tab.
+10).
+Automatic metrics.
+We employ Accuracy (Acc) and Collapsed Accuracy (CAcc)222is a relaxed accuracy wherein the choices of MCQ questions ($\geq$ 4 choices) are collapsed to become 2 choices.
+as the evaluation metrics following Hwang et al.
 (2023).
-It is worth noting that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples.
-In addition, human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$ and CoO steps of ChOiRe.
-Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via CoO.
-To this end, we randomly select $100$ answers generated by ChOiRe with ChatGPT, ChatGPT-Instruct, GPT-4, Mistral.
+Note that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples.
+Human metrics.
+Human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$, and VBN steps of COO.
+Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via VBN.
+We randomly select $100$ answers generated by COO with ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral.
 We then hire $3$ excellent undergraduates who are native English speakers as annotators.
 For FEA and LLMtop-$K$ steps, each annotator is instructed to rate on a 1-3 scale (3 is best) via the Satisfaction criterion defined as how well the algorithm of LLMs performs in filtering/ranking, subjectively.
-To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the CoO explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions.
+To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the VBN explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions.
 Three annotators are also guided to rate the criteria on a 1-3 scale.
 Each metric’s final score is the average of three annotators’ scores.
-The scoring instructions are in section D.1 and the inter-annotators’ agreement is assessed by Kripp’s alpha Krippendorff (2011).
-5 Main Results
-Model ChatGPT ChatGPT-Inst GPT-4 Mistral-7B-Ins.-v0.2
-W/o persona 46.60/65.72 44.91/63.60 - 41.24/59.54
-DIO-top8 50.22/69.21 51.95/71.16 57.98/76.86 44.16/62.47
-DIO-top8 + Self-refine 43.14/65.33 42.71/62.98 - 36.23/55.06
-DIO-top8 + CoT 49.96/69.05 51.90/71.51 - 52.25/71.95
-DIO-top8 + SC 50.58/69.66 52.06/71.87 - 53.14/72.88
-DIO-top8 + FEA 50.64/69.85 52.63/72.30 - 44.99/64.09
-DIO-top8 + CoO 50.97/70.22 52.08/71.65 - 53.79/73.59
-DIO-LLMtop8 51.03/70.31 52.80/72.60 - 45.86/64.98
-DIO-LLMtop8 + FEA 51.19/70.69 52.97/72.84 - 45.23/64.73
-DIO-LLMtop8 + FEA + CoO 51.90/71.57 53.01/72.91 59.02/78.70 54.21/74.09
-ChOiRe 52.21normal-$\dagger$/72.09normal-$\dagger$ 53.26normal-$\dagger$/73.26normal-$\dagger$ 59.30normal-$\dagger$/78.82normal-$\dagger$ 54.43normal-$\dagger$/74.34normal-$\dagger$
-% Improvements +3.22/+3.49 +2.52/+1.93 +2.28/+2.55 +2.42/+2.00
-Table 1: Overall Accuracy/Collapsed Accuracy on ChatGPT, ChatGPT-Instruct, and GPT-4.
-FEA is our first step, Filtering Explicit Attributes.
-LLMtop8 is the second step, ranking and selecting top-8 implicit persona opinions as demonstrations, CoO is Chain-of-Opinion reasoning.
-Improvements are calculated with the best baseline.
-$\dagger$ denotes our model outperforms baselines significantly with p-value < 0.01 under t-test.
-Model GPT-2-base GPT-2-large FlanT5-base FlanT5-large
-W/o persona 41.14/58.87 21.94/39.11 48.98/68.33 39.83/58.43
-DIO-top8 21.23/38.64 24.94/42.22 55.00/74.98 54.94/74.79
-DIO-top8 + FEA 22.62/40.97 25.65/45.21 55.78/75.34 58.77/77.26
-DIO-LLMtop8 22.65/41.12 28.86/47.60 57.97normal-$\dagger$/77.46normal-$\dagger$ 58.20/77.56
-DIO-LLMtop8 + FEA 25.05/44.41 29.54normal-$\dagger$/48.66normal-$\dagger$ 57.45/77.13 59.00normal-$\dagger$/78.46normal-$\dagger$
-% Imp. over DIO-top8 +17.99/+14.93 +18.44/+15.25 +5.40/+3.30 +7.38/+4.90
-Table 2: Performance of fine-tuned baselines with our proposed FEA and LLMtop8 steps preprocessed by ChatGPT.
-$\dagger$ denotes our model significantly outperforms baselines with p-value < 0.01 under t-test.
-Overall Prompting Results.
-Table 1 shows our macro experimental outcomes.
-We derive $4$ main observations in this task.
-First, ChOiRe improves the best among baselines significantly with $3.22\%$, $2.52\%$, $2.28\%$, $2.42\%$ accuracy for ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral.
-It establishes a strong SOTA result with GPT-4, surpassing previous SOTA DIO-top8 with InstructGPT achieving 53.74% Hwang et al.
-(2023) by a notable margin.
-Notably, in the case of GPT-4, we utilize ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one.
-Second, we see that Accuracy and Collapsed Accuracy have the same trend, and ChOiRe also achieves the SOTA on Collapsed Accuracy with the highest improvement of $3.94\%$ observed with ChatGPT.
-Third, naïve CoT ("answer the following question step-by-step") helps Mistral but slightly harms ChatGPT and ChatGPT-Instruct with DIO-top8 Hwang et al.
+The scoring instructions are in Section D.1 and the inter-annotators’ agreement is assessed by Krippendorff’s alpha Krippendorff (2011).
+5.2 Main Results
+We outline the prompting, fine-tuning, fine-grained, and human evaluation results of COO.
+Prompting results.
+Tab.
+2 shows our main experimental outcomes.
+For GPT-4, it attains 57.98% Acc with DIO-top8 and 59.42% with COO, establishing a strong SOTA result surpassing the previous of 53.74% Hwang et al.
 (2023).
-On the other hand, SC improves all models.
-Therefore, we attribute CoT’s limitation to the inconsistency of its explanations (3).
-Meanwhile, ChOiRe with CoO consistently attains improvements, verifying the effectiveness of explicitly requiring the model to analyze all the personae.
-Finally, ChatGPT, ChatGPT-Instruct, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average, respectively.
-This suggests that over half of explicit personae may be noisy for models to predict opinions.
-Fine-grained Prompting Results.
-Diving deeper into the benchmark topics in table 5, ChOiRe achieves SOTA results in $8/15$, $8/15$, $11/15$, $13/15$ topics for ChatGPT and ChatGPT-Instruct, GPT-4, and Mistral.
-The improvements are especially huge for some topics.
-For example, compared with the best among baselines, it improves GPT-4 up to $12.08\%$ accuracy on Views on gender, ChatGPT up to $9.82\%$ on Economic Inequality.
-We also specifically compare ChOiRe with the best baseline DIO-top8 + SC in section B.3, showing $8/12$ improvements for ChatGPT and ChatGPT-Instruct.
-We further plot the accuracy distribution over users of ChOiRe, specifically for ChatGPT in fig.
+Overall, COO delivers the best results with significant improvements of 2-4% Acc for all benchmarked LLMs with ChatGPT securing the most gain of 4.11%.
+Its component steps consistently enhance baselines with VBN > LLMtop8 > FEA on average.
+Among prompting methods, we observe that naïve CoT helps Mistral slightly but harms ChatGPT and ChatGPT-it, while SC improves all.
+We attribute this to the inconsistency and unreliability of CoT reasoning (4), and the challenge of this task.
+Meanwhile, COO’s VBN reasoning consistently improves, verifying the effectiveness of requiring LLMs to explicitly analyze all the personae through values, beliefs, and norms.
+Additionally, self-refine consistently lowers model performance, indicating that multiple refinement rounds may be counterproductive for this complex task as these rounds may amplify the model’s inherent biases Xu et al.
+(2024b), leading to more biased predictions.
+Among models, notably, for GPT-4, we use ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one.
+Finally, ChatGPT, ChatGPT-it, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average.
+This suggests that over half of explicit personae are noisy for model opinion prediction.
+Fine-tuning results.
+Tab.
+2 presents our fine-tuning results.
+Overall, leveraging COO’s FEA, LLMtop-$K$, and VBN extra variables results in significant improvements for both decoder-only and encoder-decoder models, with average gains of 22.20% for GPT-2s and 8.93% for FlanT5s.
+Among COO’s steps, FEA contributes the least, while LLMtop-$K$ and VBN extra variables drive more substantial gains across most models.
+The VBN extra variables can be seen as distilled knowledge from ChatGPT, intuitively enhancing fine-tuning results.
+However, notably, COO’s FEA and LLMtop-$K$, which focus on selecting relevant explicit and implicit personae, already deliver substantial improvements across all models, bringing FlanT5-large’s performance on par with GPT-4.
+This verifies our hypothesis that removing irrelevant personae is necessary for high performance.
+Finally, GPT-2-base performs surprisingly well even without user demographics and ideology, possibly due to potential contamination Sainz et al. (2023) with public polling data from OpinionQA.
+Fine-grained Results.
+We compare COO with the DIO-top8 baseline across 15 OpinionQA topics to assess its performance in detail.
+Overall, we see COO consistently outperforms DIO-top8 in most topics, with the largest improvements in “View on gender” (+17.69% with Mistral), “Autonomous vehicles” (+13.49 with GPT-2), and “Misinformation” (+11.61 with ChatGPT).
+These gains further emphasize COO’s effectiveness in enhancing LM performance on social and belief-driven topics, especially those involving complex reasoning.
+Full results are provided in Appx.
+Tab.
+6.
+Human evaluation results.
+Model FEA Satis. LLMtopK Satis. VBN Rea. VBN FI
+ChatGPT $2.56_{\alpha=0.74}$ $\textbf{2.32}_{\alpha=0.68}$ $2.85_{\alpha=0.88}$ $\textbf{2.91}_{\alpha=0.90}$
+ChatGPT-it $\textbf{2.64}_{\alpha=0.71}$ $2.28_{\alpha=0.65}$ $2.87_{\alpha=0.90}$ $\textbf{2.95}_{\alpha=0.87}$
+GPT-4 $\textbf{2.90}_{\alpha=0.91}$ $2.21_{\alpha=0.77}$
+Mistral $2.31_{\alpha=0.65}$ $2.12_{\alpha=0.64}$ $2.58_{\alpha=0.68}$ $2.16_{\alpha=0.55}$
+Table 3: Human evaluation results. LLMs perform adequately in COO’s FEA step, and excel in VBN reasoning, but face challenges with the LLMtop-K step. $\alpha$ denotes the Krippendorff’s alpha.
+For COO’s FEA and LLM ranking steps, from Tab.
+3, ChatGPT and ChatGPT-it generally achieve similar performance and are better than Mistral: ChatGPT excels slightly in ranking while ChatGPT-it performs slightly better in performing FEA.
+Three models are proficient in FEA but struggle with the ranking task where the common error is misplacing relevant opinions, due to this task’s complexity.
+Second, four models effectively generate VBN reasoning thoughts, and GPT-4 performs the best.
+Finally, ChatGPT and ChatGPT-it follow our instructions to explain and analyze the explicit and implicit personae provided one by one with VBN significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$.
+Our hypothesis is they are optimized for following instructions, while GPT-4 is optimized for completing texts.
+6 Discussion
+We discuss the main analyses here, including COO generalization (Section 6.1) and COO’s Steps sequentially. Extra analyses are supplemented in Appendix B.
+Method ChatGPT
+W/o personae 46.60
+COO w/o personae (Step 4 activated) 47.79
+DIO-top8 w/o explicit 49.22
+COO w/o explicit (Steps 2, 3 (PBN), 4 activated) 51.66
+DIO-top8 w/o implicit 47.16
+COO w/o implicit (Steps 1, 3 (EV), 4 activated) 50.13
+Table 4: COO’s results with missed persona(e) with ChatGPT. In all scenarios, COO outperforms the baselines significantly.
+6.1 COO in Missing Personae Scenarios
+COO demonstrates strong generalizability even when explicit, implicit, or both types of personae are missing.
+Specifically, in the absence of both, COO reduces to Self-consistency Wang et al.
+(2023b); without explicit personae, only Step 1 of COO is skipped; without implicit personae, Step 2 is omitted.
+As shown in Tab.
+4, COO consistently outperforms the leading baseline by an absolute margin of 2-3% in all scenarios.
+Finally, its steps can be generalizable to other tasks, see Section B.8.
+6.2 Method Analysis
+FEA: ablation study.
+To gauge the impact of removing irrelevant explicit personae (FEA), we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al.
+(2023), denoted as DIO-top8 + FEA in Tab.
+2.
+We observe a 1-2% Acc performance boost on ChatGPT, ChatGPT-it, and Mistral respectively.
+This underscores the effectiveness of eliminating irrelevant explicit personae in improving the model prediction.
+FEA: irrelevant personae distribution.
+To understand the explicit personae filtered by LLMs across various topics, we document the top 3 removed personae in Section B.3.
+"Citizenship" is observed to be the most frequently removed attribute, followed by "Race".
+This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation.
+Another explanation may be the lack of correlation between citizenship/race and opinions in the OpinionQA dataset.
+Additionally, we also see that ChatGPT often categorizes “Marital status" as non-useful, ChatGPT-it commonly removes “Frequency of religious attendance", and “Gender" got removed by Mistral, revealing potential biases in LLMs.
+LLMtop-$K$: compared to semantic top-$K$.
+From Tab.
+2, DIO-LLMtop8 outperforms DIO-top8 by $1-4\%$ accuracy on ChatGPT, ChatGPT-it, and Mistral, confirming that prioritizing usefulness over semantic similarity improves model prediction.
+To further understand the difference between semantic similarity orders and usefulness orders, we discuss (1) the agreement of LLM orders and semantic similarity orders, and (2) maximum disagreement points between these orders.
+In tackling (1), we calculate Kendall’s Tau coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-it, Mistrial, and semantic similarity orders, and the results are presented in Appx.
+Fig.
 4.
-We see that the majority accuracy is $0.5$, with a few users scoring zero and over $20$ achieving perfection.
-Fine-tuning Results.
-Table 2 presents our fine-tuning outcomes.
-Notably, leveraging the ChOiRe’s FEA and LLMtop-$K$ steps on the fine-tuning data yields substantial enhancements for GPT-2-large and FlanT5-large, showcasing relative accuracy improvements of $18.44\%$ and $7.38\%$ respectively.
-Remarkably, ChOiRe’s FEA and LLMtop-$K$ steps bring FlanT5-large’s performance on par with GPT-4, despite GPT-4’s significantly stronger capability.
-Furthermore, ChOiRe’s LLMtop-$K$ proves particularly beneficial for enhancing FlanT5-base.
-Surprisingly, GPT-2-base performs well even without user demographic and ideological information, possibly due to potential contamination Sainz et al.
-(2023) with public polling data from OpinionQA.
-Human Evaluation Results.
-Model FEA Satis. LLMtopK Satis. Rea. Foll. Inst.
-[…]
-[…]
-[…]
-[…]
-[…]
-Our human evaluation results in table 3 reveal three key findings.
-First, ChatGPT and ChatGPT-Instruct achieve similar performance in filtering explicit personae and ranking opinions, while Mistral achieves lower results.
-While ChatGPT excels slightly in ranking, ChatGPT-Instruct performs slightly better in explicit personae selection.
-Three models proficiently filter unnecessary explicit personae, but ranking opinions poses a more challenging task intuitively and empirically, with a common error being the inconsistent relevance ranking of opinions, sometimes misplacing high-level relevance.
-Second, four models effectively generate reasonable thoughts leading to the final answer, and GPT-4 performs the best.
-Finally, ChatGPT and ChatGPT-Instruct follow our instructions to explain and analyze the explicit and implicit personae provided one by one with CoO significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$.
-We hypothesize that this is because ChatGPT and ChatGPT-Instruct excel in following instructions, while GPT-4 is optimized for completing texts.
-6 Discussion
-We discuss the main analyses in this section. Extra analyses are presented in appendix B.
-6.1 Methodology Analysis
-Ablation of FEA.
-To gauge the impact of filtering unnecessary explicit personae (FEA) on performance, we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al.
-(2023), denoted as DIO-top8 + FEA in table 1.
-The results indicate enhancements with DIO-top8 + FEA achieving a 0.8%, 1.3%, 1.9% accuracy performance boost on ChatGPT, ChatGPT-Instruct, and Mistral respectively.
-This underscores the effectiveness of eliminating irrelevant explicit personae in improving the models’ ability to understand and predict human opinions.
-FEA via Topics.
-To understand the explicit personae filtered by LLMs across various topics, we document the top-3 removed personae in section B.5.
-We observe that "Citizenship" is consistently the most frequently removed attribute, followed by "Race".
-This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation.
-Another explanation may be the lack of correlation between citizenship/race and opinions in the US-centric OpinionQA dataset.
-Additionally, we also see that ChatGPT often categorizes ‘‘Marital status" as non-useful, ChatGPT-Instruct commonly removes ‘‘Frequency of religious attendance", and ‘‘Gender" got removed by Mistral, revealing potential biases in LLMs.
-LLMtop-$K$ versus Top-$K$.
-From table 1, DIO-LLMtop8 outperforms DIO-top8 by $1.6\%$, $1.6\%$, $3.8\%$ accuracy on ChatGPT, ChatGPT-Instruct, Mistral confirming that prioritizing meaning and usefulness improves opinion prediction.
-One possible explanation for this can be the orders ranked by semantic similarity scores only consider ranking with the input questions Hwang et al.
-(2023), while our orders consider both input questions and their answer choices (fig.
-1).
-We further explore two key aspects: (1) The agreement of LLM-orders and semantic similarity orders, and (2) Points of maximum disagreement between these orders.
-To measure the ranking agreements, we calculate Kendall’s Tau correlation coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-Instruct, and Mistrial and orders sorted by semantic similarity scores, and the results are presented in fig.
-6 and fig.
-7.
-Surprisingly, for ChatGPT and ChatGPT-Instruct, we find that the two ranking orders have minimal monotonous relations with means approximating 00 and low standard deviations showing no agreement.
-For Mistral, we find a low agreement with a mean of $0.43$ score.
+Surprisingly, for ChatGPT and ChatGPT-it, we find that the two ranking orders have no agreement with means approximating 00.
+For Mistral, we observe a low agreement with a mean of $0.43$ score.
 These low and no agreements further verify that ranking by usefulness can be very different from ranking by semantic similarity.
-We also deep dive into cases with notable order variations to address (2).
-Section E.2 illustrates one such case in the "Guns" topic.
-We derive three observations.
-First, not all top-8 opinions by semantic similarity scores help predict the opinion.
+For answering (2), Section E.2 illustrates one such case in the "Guns" topic.
+We observe that not all top-8 opinions by semantic similarity scores help predict the opinion.
 For example, the $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question.
@@
 The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership.
-Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
-Opinions Order Analysis of LLMtop-$K$ Step.
-The performance difference of DIO-top8 and DIO-LLMtop8 in table 1 highlights that LLMs are sensitive to the chosen implicit personae opinions.
-An important question arises: Are LLMs also affected by the input order of implicit persona opinions in the ranking step (section 3)?
-Our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large.
+Finally, by using the LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
+LLMtop-$K$: the order of input opinions.
+We study the (1) sensitivity and the (2) performance variance of LLMs to the order of input implicit personae in the LLMtop-$K$ step.
+To address (1), our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large ($K\geq 8$).
 We randomly select $300$ questions, shuffle implicit persona opinions four times with different seeds, and record four LLM ranking outputs for each.
 We also collect one more LLM ranking output by feeding implicit personae opinions in semantic similarity order.
-For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap coefficient Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$.
-The scores, shown in fig.
-5, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models.
-Despite this, is there substantial variance in model performance across random seeds?
-Our findings reveal no significant variance, with the variants statistically outperforming the baseline DIO-top8.
+For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap Coefficient (OC) Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$.
+The scores, shown in Appx.
+Fig.
+3, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models.
+For (2), we find no significant performance variance.
 Specifically, we assess ChatGPT and Mistral with DIO-LLMtop8 on $3$ out of $4$ random seeds, detailed in Section B.5.
 The results demonstrate relatively small standard deviations in their performance, and critical values of $99\%$ CI of DIO-LLMtop8 under t-test for both models surpass DIO-top8, confirming that LLMtop8’s effectiveness is not due to randomness.
-CoO versus CoT.
-Figure 2: Consistency scores of the baseline DIO-top8 (ChatGPT) with CoO and CoT.
-Table 1 indicates that Chain-of-Thought (CoT) Kojima et al.
-(2022) slightly harms baseline DIO-top8 performance for ChatGPT and ChatGPT-Instruct.
-Conversely, our Chain-of-Opinion reasoning (CoO) enhances overall performance for all models.
-To investigate the consistency of CoT and CoO, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and CoO, at $3$ different temperatures $0.3$, $0.6$, $0.9$.
-For each prompting technique, we measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score.
-The results are illustrated in fig.
-2 showing that CoO brings slightly better consistent answers compared to CoT, especially when the temperature is high which verifies CoO potentially enhances the reliability of LLMs.
-Dynamic Numbers of Opinions Analysis.
-Table 4 illustrates our analysis answering two research questions: (1) How frequent can’t LLMs answer the question?
-and (2) How do LLMs perform when more opinions than $K=8$ are provided in ChOiRe?
-Our findings show that, firstly, with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions.
-Secondly, increasing the number of opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in section 3.
-Lastly, while including more opinions could possibly harm the performance of models, our answer consistency strategy enables LLMs to achieve the best results across three different K values.
-Model ChatGPT ChatGPT-Inst GPT-4 Mistral
-% of ITA of DIO-LLMtop8 + FEA + CoO 0.61 1.32 9.71 0.00
-DIO-LLMtop8 + FEA + CoO 51.90 53.01 59.02 54.21
-% of ITA of DIO-LLMtop10 + FEA + CoO 0.12 1.01 5.44 0.00
-DIO-LLMtop10 + FEA + CoO 51.55 52.74 58.88 53.88
-% of ITA of DIO-LLMtop12 + FEA + CoO 0.00 0.66 3.12 0.00
-DIO-LLMtop12 + FEA + CoO 51.60 52.31 59.11 52.96
-ChOiRe 52.21 53.26 59.30 54.43
-Table 4: Extra analysis on ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. ITA stands for "Impossible To Answer".
-6.2 Error Analysis
-FEA Misses Key Explicit Personae.
-Despite showing promising results in removing unuseful explicit personae depicted in table 3, we observe that LLMs sometimes misselect relevant personae.
-One such example is the top-left of section E.6.
-We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
-LLMtop-$K$ Opinions Include Less Relevant Ones.
-While LLMs generally demonstrate a commendable ability to rank implicit opinions by usefulness, as exemplified in section E.2, we also observe they frequently include less relevant, or even irrelevant opinions to the ranked list such as in section E.6-bottom.
-We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
-LLMs May Not Follow the Instructions.
-Although ChatGPT and ChatGPT-Instruct demonstrate a robust ability to adhere to our instructions for opinion prediction via CoO, the same level of proficiency is not observed in Mistral and GPT-4, as shown in section E.6-top-right.
-We posit this disparity arises from the fact that ChatGPT and ChatGPT-Instruct excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.
+VBN: compared to CoT.
+Tab.
+2 indicates that Chain-of-Thought (CoT) Kojima et al.
+(2022) slightly harms the performance for ChatGPT and ChatGPT-it.
+Conversely, our Value-Belief-Norm (VBN) reasoning enhances performance for all models.
+To investigate the consistency of CoT and VBN, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and VBN, at $3$ different temperatures $0.3$, $0.6$, $0.9$.
+We measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score.
+The results are illustrated in Appx.
+Fig.
+2 showing that VBN brings better consistent answers compared to CoT, especially when the temperature is high verifying VBN potentially enhances the reliability of LLMs.
+Answer consistency with dynamic opinions.
+We study (1) how frequently LLMs are unable to answer the question and (2) the impact on performance when more than $K=8$ opinions are provided.
+Tab.
+5 provides the results.
+We find that with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions.
+Increasing the #opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in Section 4.
+Lastly, while increasing $K$ could harm the model performance, COO’s answer consistency enables LLMs to achieve the best results across K values.
+Model ChatGPT ChatGPT-it GPT-4 Mistral
+ITA of DIO-LLMtop8 + FEA + VBN 0.20 0.91 3.40 0.00
+DIO-LLMtop8 + FEA + VBN 52.16 53.08 59.11 54.56
+ITA of DIO-LLMtop10 + FEA + VBN 0.00 0.00 1.43 0.00
+DIO-LLMtop10 + FEA + VBN 51.89 52.90 58.88 53.62
+ITA of DIO-LLMtop12 + FEA + VBN 0.00 0.00 0.00 0.00
+DIO-LLMtop12 + FEA + VBN 51.60 52.03 59.18 54.21
+COO 52.66 53.58 59.42 54.40
+Table 5: Percentage of “Impossible To Answer” (ITA) (%) observed with corresponding performance during generation.
 7 Conclusion
-We propose ChOiRe, a four-step solution framework for individual opinion prediction via differentiating the utilization of user’s explicit versus implicit personae.
-We further introduce Chain-of-opinion reasoning and answer consistency over variable numbers of input implicit personae guiding the models to derive thorough predictions.
-ChOiRe achieves strong SOTA results with limited inference calls, demonstrating its strong effectiveness.
-Additionally, Steps (i) and (ii) of ChOiRe significantly improve the fine-tuning of opinion-aligned models.
-We strongly suggest that our method should only be used for positive moral intents, avoiding making LLMs echo chambers Vicario et al.
-(2016).
-In the future, we will focus on developing frameworks that utilize explicit and implicit personae more efficiently.
+This paper identifies two major challenges in aligning LLMs with human opinions via personae: noisy personae and ineffective reasoning strategies over personae.
+To address these, we propose COO, a novel four-step framework in the light of the Value-Belief-Norm theory: the first two steps tackle the noise issue using LLMs as analysts, while the last two enhance reasoning through the novel VBN reasoning.
+COO significantly improves LLM prompting and fine-tuning, demonstrating high generalizability in scenarios with missing personae and potential applications in other personalization tasks.
+Acknowledgements
+This research project is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL), the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207), and the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005).
+Do Xuan Long is supported by the A*STAR Computing and Information Science (ACIS) Scholarship.
 Limitations
-One limitation of our proposed ChOiRe framework is that it requires the LLMs to have a good capability in following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions one by one via CoO.
+One limitation of COO is that it requires the LLMs to be well capable of following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions VBN reasoning.
 However, we foresee that this limitation is going to be overcome by cutting-edge AI language models, in the present and near future.
-Additionally, our method also utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world.
-However, to what extent is the personal information provided, our ChOiRe is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions.
+Additionally, COO utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world.
+However, to what extent is the personal information provided, our COO is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions (see Section 6.1).
+Finally, for fine-tuning, COO currently leverages ChatGPT for data synthesis, which may pose challenges to replicability.
+To address this, we will open-source our generated data and code here.
+Future research could explore using open-source LLMs, which are increasingly powerful and comparable to proprietary models.
 Ethical Considerations
@@
 Bias amplification and fairness.
-A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Vicario et al.
+A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Del Vicario et al.
 (2016).
 This can ultimately harm users by reinforcing polarized or undesirable views.
-To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed ChOiRe involves presenting user demography or ideology group responses alongside personalized answers.
+To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed COO involves presenting user demography or ideology group responses alongside personalized answers.
 Additionally, COO can encourage users to reflect on their previous viewpoints.
@@
 This customization can help mitigate potential biases and provide individuals with AI-generated opinions that align more closely with their values and preferences.
+Using ChatGPT for data synthesis.
+We follow prior studies (Ray, 2023; Tan et al., 2024) to use OpenAI ChatGPT to synthesize data.
+Additionally, we obey OpenAI’s terms of use [^3] https://openai.com/policies/row-terms-of-use/ to use ChatGPT’s synthesized data to develop models that do not compete with OpenAI.
+Misuse and responsibility for long-term societal impact.
+While our method aims to align opinions with individuals, it also introduces risks of misuse, such as the propagation of harmful ideologies or manipulating human opinions.
+While this is not what our method is designed for, there is no way to prevent this type of misuse.
+We emphasize the importance of ethical alignment in deploying these systems and suggest that developers and practitioners must establish robust guidelines and oversight mechanisms to prevent misuse.
+Furthermore, incorporating mechanisms to monitor and audit AI-generated content following ethical norms is crucial.
+We also recognize potential risks of unintended societal consequences, such as fostering group biases or undermining collective decision-making processes.
+We also acknowledge the potential for unintended societal consequences, such as fostering group biases or undermining collective decision-making processes.
+To mitigate these risks, we recommend integrating ethical safeguards, including mechanisms to monitor and audit AI use, and fostering user awareness and diverse perspectives during model training and deployment.
+This approach aims to minimize negative societal impacts and ensure the technology is applied constructively and equitably.
 Human evaluation.
 Through human evaluations, we observe that our proposed method does not generate any discriminatory, insulting responses.
-We validate the intermediate steps of our proposed ChOiRe by human evaluation which involves manual labor.
-We hire annotators to score, and the hourly pay is set to \$$\$15$, which is higher than the local statutory minimum wage.
+We validate the intermediate steps of our proposed COO by human evaluation which involves manual labor.
+We hire annotators to score, and the hourly pay is set to \$$\$20$, which is higher than the local statutory minimum wage.
 Therefore, we do not anticipate any major ethical concerns raising from human evaluations.
-Appendix A Baselines
-A.1 Baselines Implementation Details
+Appendix A Implementation Details
 Prompting.
 […]
-Model ChatGPT/ChatGPT-Inst/GPT-4/Mistral
-Guns Auto. & driverless vehicles Views on gender Com. types & sex. harassment Race
-W/o persona 53.07/37.30/——-/30.48 47.73/48.26/——-/41.72 50.53/42.94/——-/37.39 47.73/41.67/——-/29.34 41.95/45.28/——-/37.55
-DIO-top8 53.87/57.00/60.39/44.73 45.33/44.78/53.22/41.72 53.21/52.15/63.73/40.09 43.47/45.24/42.86/35.45 43.06/44.65/55.17/41.11
-DIO-top8 + CoT 54.55/52.33/——-/55.48 47.22/46.77/——-/49.00 48.11/57.67/——-/54.28 42.39/42.26/——-/42.01 45.63/43.40/——-/49.78
-DIO-top8 + SC 54.40/52.85/——-/56.57 43.73/48.26/——-/52.31 55.61/56.44/——-/56.30 45.33/40.48/——-/42.01 45.00/43.40/——-/50.00
-ChOiRe 57.06/58.21/63.37/58.00 49.25/51.92/50.00/53.75 59.23/53.07/71.43/57.78 39.88/44.14/47.96/42.08 42.77/47.28/50.57/51.44
-Gender & Leadership America in 2050 Trust in science Biomedical & food issues Misinformation
-W/o persona 53.13/50.83/——-/43.51 39.73/39.13/——-/41.95 50.40/47.29/——-/48.34 53.87/53.63/——-/53.21 46.93/40.38/——-/53.63
-DIO-top8 48.27/54.70/65.55/50.23 46.93/46.20/43.70/35.14 54.93/61.58/61.54/51.65 52.27/55.86/58.03/52.78 49.33/52.11/52.71/50.77
-DIO-top8 + CoT 48.58/50.83/——-/55.79 43.05/48.91/——-/43.76 54.10/65.02/——-/58.28 56.91/57.54/——-/57.08 49.57/53.99/——-/53.19
-DIO-top8 + SC 49.07/53.60/——-/57.87 45.87/47.83/——-/46.03 56.27/65.52/——-/58.94 53.07/57.54/——-/58.58 45.00/53.52/——-/53.85
-ChOiRe 52.22/57.78/63.03/57.87 49.46/48.99/45.37/47.50 56.43/55.50/68.46/60.37 54.75/57.26/61.61/58.58 46.45/53.62/57.36/53.85
-Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views
-W/o persona 43.24/40.28/——-/33.64 47.06/44.36/——-/46.08 43.67/49.15/——-/34.07 46.13/46.71/——-/40.42 40.80/48.95/——-/46.20
-DIO-top8 53.24/47.22/47.73/43.31 57.22/57.89/62.50/47.42 45.60/51.98/63.81/41.87 49.60/57.23/66.67/41.27 56.80/46.85/62.07/44.13
-DIO-top8 + CoT 53.38/47.22/——-/56.91 59.57/55.64/——-/54.36 47.65/51.98/——-/51.45 46.42/56.58/——-/51.06 53.30/45.45/——-/50.80
-DIO-top8 + SC 54.05/47.22/——-/58.06 55.35/54.89/——-/57.04 46.13/51.98/——-/52.76 46.42/55.26/——-/52.89 57.33/47.55/——-/51.67
-ChOiRe 54.29/53.33/52.27/58.06 60.00/58.77/63.89/58.50 52.33/50.13/64.76/51.89 44.74/55.26/64.58/52.76 51.05/53.74/67.82/53.34
-Table 5: Fine-grained accuracy results of ChatGPT/ChatGPT-Instruct/GPT-4/Mistral. DIO stands for Demographic + Ideology + Opinions (section 4).
 Fine-tuning.
 […]
-Appendix B Extra Analysis
-B.1 Additional Baseline Comparisons
+Appendix B Additional Analyses
+B.1 Fine-grained Analyses
+Guns Auto. vehicles Views on gender Community types Race
+ChatGPT 53.87 / 57.06 45.33 / 51.25 53.21 / 59.23 43.47 / 40.88 43.06 / 43.27
+ChatGPT-it 57.00 / 58.21 44.78 / 51.92 52.15 / 53.07 45.24 / 46.14 44.65 / 48.28
+Mistral 44.73 / 58.12 41.72 / 53.75 40.09 / 57.78 35.45 / 42.08 41.11 / 51.44
+GPT-4 60.39 / 63.37 53.22 / 50.00 63.73 / 71.43 42.86 / 47.96 55.17 / 50.57
+[…]
+GPT-2 large 25.34 / 30.07 26.04 / 31.96 22.66 / 24.66 23.33 / 29.00 21.67 / 22.42
+FlanT5 62.08 / 60.39 55.60 / 59.61 57.45 / 62.60 45.98 / 47.20 54.23 / 61.56
+FlanT5 large 60.56 / 65.45 51.52 / 58.87 60.28 / 60.64 46.21 / 49.76 49.24 / 55.82
+Gender & Leadership America in 2050 Trust in science Bio. & food issues Misinformation
+ChatGPT 48.27 / 52.22 46.93 / 49.46 54.93 / 56.43 52.27 / 54.75 49.33 / 60.94
+ChatGPT-it 54.70 / 56.28 46.20 / 49.00 61.58 / 55.50 55.86 / 57.26 52.11 / 53.62
+Mistral 50.23 / 57.87 35.14 / 47.60 51.65 / 60.37 52.78 / 58.58 50.77 / 53.85
+GPT-4 65.55 / 63.03 53.71 / 45.27 61.54 / 68.46 58.03 / 61.61 52.71 / 57.26
+[…]
+GPT-2 large 21.50 / 28.11 27.90 / 29.27 29.61 / 30.96 28.60 / 33.52 23.34 / 30.13
+FlanT5 63.98 / 66.08 55.82 / 55.00 64.22 / 63.00 61.41 / 61.56 60.49 / 60.00
+FlanT5 large 58.54 / 64.00 46.54 / 51.30 61.76 / 68.43 57.34 / 63.75 51.77 / 61.97
+Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views Average
+ChatGPT 53.24 / 54.29 57.22 / 61.00 45.60 / 52.43 49.60 / 45.54 56.97 / 51.15 50.22 / 52.66
+ChatGPT-it 51.02 / 54.33 57.89 / 58.25 51.98 / 50.13 57.23 / 57.86 46.85 / 53.84 51.95 / 53.58
+Mistral 43.31 / 58.06 47.42 / 58.50 41.87 / 51.89 42.0 / 52.76 44.13 / 53.34 44.16 / 54.40
+GPT-4 47.73 / 52.27 62.50 / 63.89 63.81 / 64.76 66.67 / 63.58 62.07 / 67.82 57.98 / 59.42
+[…]
+GPT-2 large 29.53 / 37.30 24.19 / 28.57 22.28 / 28.21 24.87 / 32.01 23.38 / 28.46 24.94 / 30.21
+FlanT5 58.36 / 58.00 64.59 / 65.60 54.06 / 57.60 63.23 / 58.04 49.08 / 52.80 55.00 / 59.62
+FlanT5 large 57.92 / 61.53 64.98 / 67.02 51.66 / 58.72 57.19 / 57.95 48.55 / 56.67 54.94 / 60.13
+Table 6: Fine-grained accuracy results of models with DIO-top8 Hwang et al. (2023) / Chain-of-Opinion (COO; ours).
+Tab.
+6 presents our fine-grained results across topics of the OpinionQA dataset Santurkar et al.
+(2023).
+Overall, COO consistently outperforms DIO-top8 across most of the OpinionQA topics.
+The highest absolute improvements occur in “View on gender” (+17.69 with Mistral), “Autonomous vehicles” (+13.49 with GPT-2), and “Misinformation” (+11.61 with ChatGPT).
+These improvements highlight the COO’s strength in helping models better handle complex, belief-based topics, particularly those involving social and political biases.
+Across models, COO notably enhances performance, whether in strong models like GPT-4 or weaker ones like GPT-2.
+GPT-4, one of the leading LLMs for reasoning, shows marked improvement in scientific and social topics such as “Views on gender” (+7.70) and “Misinformation” (+4.55).
+GPT-2, usually less effective, also benefits from COO’s enhancements, with significant gains in “Views on gender” (+11.37) and “Political views” (+7.98).
+This further emphasizes the strength of COO in enhancing both high-performing and less capable models.
+B.2 Additional Baselines
 Model ChatGPT Mistral-7B-Instruct-v0.2
@@
 Table 7: Accuracy results of ChatGPT and Mistral with two trivial variants with two different random seeds 2000 and 2024 in Section B.2.
-In this section, we compare ChOiRe’s FEA and LLMtop-$K$ steps with two simple variants outlined in table 6.
-Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs?
+We compare COO’s FEA and LLMtop-$K$ steps with two simple variants outlined in Tab.
+7.
+Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: (1) Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs?.
+Our answer is no.
 The first variant, DIO-top8 + Random FEA, involves randomly selecting $5/12$ and $9/12$ explicit persona attributes.
-The second variant entails randomly selecting $8$ implicit persona opinions instead of using ChOiRe’s LLMtop-$K$ step.
-From table 6, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes.
+The second variant entails randomly selecting $8$ implicit persona opinions instead of using COO’s LLMtop-$K$ step.
+From Tab.
+7, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes.
 Additionally, randomly selecting $8$ implicit persona opinions also adversely affects the models, particularly ChatGPT.
-These observations underscore the effectiveness and importance of ChOiRe’s FEA and LLMtop-$K$ steps.
-B.2 Fine-grained Results of ChatGPT, ChatGPT-Instruct, and GPT-4
-Table 5 presents the fine-grained results of ChOiRe and baselines for ChatGPT, ChatGPT-Instruct, and GPT-4.
-B.3 ChOiRe versus Self-Consistency
-Fig.
-3 presents the improvements per topic comparision between ChOiRe and the SOTA baseline with Self-Consistency Wang et al.
-(2023b) DIO-top8 + SC.
-We observe that ChOiRe improves $11/15$ topics for both ChatGPT and ChatGPT-Instruct.
-Figure 3: % of improvements over the SOTA method (DIO-top8 + SC) with ChatGPT-Instruct (left) and ChatGPT (right).
-B.4 Accuracy Distribution over Users
-Fig. 4 shows the accuracy distribution over users of ChOiRe with ChatGPT. We see that the peak accuracy is at 0.5 for the majority, with a few users scoring zero and over 20 achieving perfection.
-Figure 4: Frequency distribution of accuracy over users by ChOiRe.
-B.5 Top-3 Removed Explicit Personae Attributes
-Table 7 presents the top-3 explicit personae that got removed the most by the LLMs. Among the removed personae, "Citizenship" appears to be the highest-frequency one across models, followed by "Race".
+These observations underscore the effectiveness and importance of COO’s FEA and LLMtop-$K$ steps.
+B.3 Top-3 Removed Explicit Personae Attributes
+Tab.
+8 reveals a relatively common pattern in how three LLMs, ChatGPT, ChatGPT-it, and Mistral-7B-Instruct-v0.2, filter explicit personae across various topics.
+"Citizenship" and "Race" consistently emerge as the most frequently removed attributes, suggesting a deliberate effort by these models to minimize potential biases associated with these demographic factors.
+Additionally, ChatGPT-it’s tendency to remove "Frequency" and Mistral-7B-Instruct’s broader removal of "Education" and "Political Party" highlight model-specific strategies and comprehension in filtering personae based on topic relevance.
+Overall, these patterns suggest that not all explicit personae are equally relevant for predicting opinions.
 Topic ChatGPT ChatGPT-Instruct Mistral-7B-Instruct-v0.2
@@
 Table 8: Top-3 explicit personae that got removed the most by the LLMs.
-B.6 Ranking Consistency for LLMtop-$K$ Step
-We record the average Overlap coefficient Vijaymeena and Kavitha (2016) among $5$ ranking outputs from $5$ input strategies in fig.
-5.
-The performance of those input strategies is further presented in table 8 on $300$ random samples.
+Figure 2: Consistency scores of the baseline DIO-top8 (ChatGPT) with CoO and CoT.
+B.4 Consistency and Reliability of VBN versus CoT
+Fig.
+2 presents the consistency scores of the baseline DIO-top8 (ChatGPT) with VBN and CoT reasoning over 100 samples.
+Both methods show improved consistency as temperature increases, with scores rising from about 58%-82% for CoT, and 62%-88% for VBN.
+VBN consistently outperforms CoT across temperatures, suggesting that it is more robust and reliable compared to CoT.
+B.5 Ranking Consistency for LLMtop-$K$ Step
+We compute the average pairwise Overlap coefficient Vijaymeena and Kavitha (2016) $OC(A,B)=\frac{|A\cap B|}{\min(|A|,|B|)}$ across five ranking outputs generated by five input strategies Fig.
+3.
+The performance of these strategies, evaluated on 300 random samples, is detailed in Tab.
+9.
+Results show a negligible variance across three different random seeds, indicating that randomizing the order of implicit personae for the $LLMtop8$ step yields a relatively stable strategy.
 Figure 3: ChatGPT and Mistral-7B-Instruct-v.02 overlap coefficient values for different values of $K$.
 We observe that for $K$ is large enough ($K\geq 8$), the coefficient value is relatively acceptable ($\geq 0.6$).
-Model Method Semantic Similarity Order Seed = 2024 Seed = 5 Seed = 2000 Seed = 15 Std
-ChatGPT DIO-LLMtop8 - 51.03 50.95 51.11 - 0.0652
-Mistral-7B-Instruct-v0.2 DIO-LLMtop8 - 45.86 45.55 45.36 - 0.2060
+Model Method Seed = 2024 Seed = 5 Seed = 2000 Std
+ChatGPT DIO-LLMtop8 51.03 50.95 51.11 0.0652
+Mistral DIO-LLMtop8 45.86 45.55 45.36 0.2060
 Table 9: Accuracy results of ChatGPT and Mistral on our test set with DIO-LLMtop8 where different orders of input implicit persona opinions are tested for LLMtop-K step.
-B.7 Kendall’s Tau Scores for Ranking Agreements
-Figure 6: Left: Ranking agreements between ChatGPT top-$K$ and semantic similarity top-$K$.
-Right: Between ChatGPT-Instruct top-$K$ and semantic similarity top-$K$.
+B.6 Kendall’s Tau Scores for Ranking Agreements
+Figure 4: Left / Middle / Right: Ranking agreements between ChatGPT top-$K$ / ChatGPT-it / Mistral orders and semantic similarity orders.
 One example that has a high disagreement score is shown in Section E.2.
-Figure 7: Ranking agreements between Mistral top-$K$ and semantic similarity top-$K$.
 Fig.
-6 shows our ranking agreement scores between ChatGPT and Semantic similarity metric (Left), and ChatGPT-Instruct and Semantic similarity metric (Right).
-We observe that the two ranking orders have minimal monotonous relations with means approximating 0 and low standard deviations.
+4 shows our ranking agreement between ChatGPT, ChatGPT-it, Mistral orders and semantic similarity orders.
+We observe that ChatGPT and ChatGPT-it orders have minimal monotonous relations with means approximating 0 and low standard deviations with semantic orders.
 More specifically, with ChatGPT, the maximum agreement is 0.6000 while the minimum is -0.5895 and the Kurtosis is -0.2173.
-For ChatGPT-Instruct, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017.
-B.8 Consistency Scores
-Table 9 presents the exact consistency scores for the fig.
-2.
-Besides CoO consistently outperforming CoT, we also observe that when the temperature is increased, the consistency score is decreased, which is intuitive.
-Model Temperature Consistency Score (%)
-DIO-top8 + CoT 0.3 84
-DIO-top8 + CoO 0.3 86
-DIO-top8 + CoT 0.6 79
-DIO-top8 + CoO 0.6 82
-DIO-top8 + CoT 0.9 58
-DIO-top8 + CoO 0.9 60
-Table 9: Consistency scores of CoT and CoO on $100$ random question-answer pairs. We sample $5$ answers per question and measure the % of questions that have all $5$ identical answers.
-Appendix C Prompts and Prompts Analysis
-C.1 Prompt Templates for Filtering Explicit Personae
-We present the prompt template for selecting relevant explicit personae for answering the question below.
-The template is hand-crafted and we use Chain-of-Thought (CoT) prompting Kojima et al.
-(2022) via adding "answer the above question step by step".
+For ChatGPT-it, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017.
+Meanwhile, Mistral shows a low correlation of 0.43.
+These low and no correlations highlight that usefulness orders can significantly differ from the semantic similarity orders commonly used in previous studies.
+B.7 Student T-test Results for Tab. 2
+Model Accuracy Collapsed Accuracy
+ChatGPT 4.11e-11 6.06e-13
+ChatGPT-Inst. 9.97e-8 4.45e-5
+GPT-4 4.23e-6 1.17e-9
+Mistral 6.01e-8 4.12e-6
+GPT-2-base 2.19e-69 1.82e-43
+GPT-2-large 5.62e-73 6.09e-49
+FlanT5-base 1.23e-19 3.19e-12
+FlanT5-large 2.55e-21 1.20e-17
+Table 10: The p-value computed by student t-test. We observe that all the values are significantly smaller than 0.01 verifying the significance of our improvements.
+We employ the Student t-test to assess the statistical significance between COO and the best-performing baseline for each model in Tab. 2. Essentially, under the null hypothesis:
+  • H0: There is no significant difference.
+  • H1: There is a significant difference.
+As we can see, the p-values from the tests in Tab. 10 are significantly below 0.01, indicating the significance of COO improvements.
+B.8 COO’s Generalization to Other Tasks
+Each step of COO aligns closely with methodologies from prior studies that have proven effective in personalized tasks. We outline them below:
+  • Step 1: FEA.
+Filtering irrelevant user attributes to improve generation outcomes is well-established Xu et al.
+(2024a).
+For instance, Rao et al.
+(2011); Sakaki et al.
+(2014) filter gender information using classifiers, while Kim et al.
+(2017) focus on age, and Demszky et al.
+(2019) analyze political polarity.
+Our FEA step is thus generalizable and applicable to these tasks.
+  • Step 2: LLMtop-K.
+Selecting the top-K most relevant historical opinions for the next prediction is conceptually related to re-ranking recommendations using LLMs Hou et al.
+(2024); Xu et al.
+(2024a) and selecting key utterances in dialogue generation Do et al.
+(2022).
+Our LLMtop-K step can likewise enhance personal chat and recommendation tasks.
+  • Step 3: VBN reasoning.
+The Value-Belif-Norm reasoning is an adaptation of Chain-of-Thought prompting Wei et al.
+(2022); Kojima et al.
+(2022).
+This strategy is designed for opinion prediction but can be used for tasks where user values, beliefs, and norms are crucial features.
+  • Step 4: Majority voting with a dynamic number of historical opinions. This step applies to any task where leveraging dynamic demonstrations enhances prediction accuracy.
+In conclusion, COO’s steps are well-supported by prior studies and can be generalized to benefit personalized tasks.
+Appendix C Prompts
+C.1 Cost Analyses for API Models
+Prompting costs for API models are detailed in Tab.
+11.
+For GPT-4, COO is priced similarly to the baseline DIO-top8, while DIO-top8 + SC costs nearly twice as much.
+This is because we execute the FEA and LLMtop-$K$ steps of COO using ChatGPT, which is comparatively inexpensive.
+For ChatGPT and ChatGPT-it, COO incurs an additional $7$ to $10$ US dollars compared to DIO-top8 + SC.
+However, this extra cost is justified by the significant performance gains, with particularly large improvements in certain topics.
+DIO-top8 DIO-top8 + CoT DIO-top8 + SC COO Model
+Avg. #tokens 562.72 623.62 995.89 3227.18 ChatGPT
+Total US\$ 3.01 3.73 6.82 14.05 ChatGPT
+Avg. #tokens 562.72 630.58 1019.31 3418.72 ChatGPT-it
+Total US\$ 3.12 3.84 7.11 20.11 ChatGPT-it
+Avg. #tokens 559.27 - 1021.14* 3292.66 GPT-4
+Total US\$ 91.19 - 226.15* 125.60 GPT-4
+Table 11: Prompting cost analysis of COO and other baselines as of 1st Sep 2024. * denotes our estimation on $50$ samples.
+C.2 Prompt Template for COO’s FEA
 A person can be described by the following attributes: {original_attribute_list} Based on the above list of demographic information above, now I give you a new question with possible answer choices: Question: ’{test_question}’ Answer choices: ’{test_choices}’ Please analyze which attributes in the demographic information are useful for you to answer the above question step by step.
@@
 Answer: [...]
-C.2 Prompt Templates for Implicit Feature Ranking
-We provide our hand-crafted prompt template for ranking implicit personae opinions in the usefulness order below:
+C.3 Prompt Template for COO’s LLMtop-K
 Given social behavior question-answer pairs answered by a user about his/her opinions about {subtopic}: {original_persona_question_order} You are an expert in analyzing the social behaviors of a user.
@@
 Give me the answer in the form of a Python list of indexes: Answer: [...]
-C.3 Prompt Templates for Baselines Techniques
-We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
-C.3.1 W/o Persona Santurkar et al. (2023)
-The W/o Persona prompt is provided below.
+C.4 Prompt Template for COO’s VBN Reasoning
+A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}.
+Opinions: {implicit_persona_str} Given the following question: Question: {question} Answer choices: {choice} Answer the above question by following the steps below: Analyze the user’s demographics and ideology one by one to infer their social and environmental values.
+Wrap this analysis by <EV> and </EV>.
+Analyze the user’s historical opinions to infer their beliefs and norms from their social and environmental values.
+Wrap this analysis by <PBN> and </PBN>.
+From the above analyses, which opinion he is likely to choose?
+Answer: A.
+or B.
+or C.
+or D.
+or E....
+C.5 Prompt Templates for Baselines
+We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-it OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
+C.5.1 W/o Persona Santurkar et al. (2023)
 Question: {question} Answer choices: {choice} Complete the answer by the following format without any explanation: Answer: A. or B. or C. or D. or E...
-C.3.2 DIO-top8 Hwang et al. (2023)
-The DIO-top8 prompt is provided below.
+C.5.2 DIO-top8 Hwang et al. (2023)
 A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}.
@@
 or E....
-C.3.3 Self-refine Madaan et al. (2023)
-The Self-refine prompts Madaan et al. (2023) are provided below, feedback step and refine step respectively.
+C.5.3 Self-refine Madaan et al. (2023)
 You are given a question and an answer for that question.
@@
 Question: {test_question} Choices: {choices} Answer: {selected_choice} Feedback:
-You are given a question, an answer to that question and a feedback to the answer.
+You are given a question, an answer to that question, and feedback to the answer.
 Based on the feedback, refine your answer and generate the final answer in around 170 words.
 Question: {test_question} Answer: {selected_choice} Feedback: {feedback} Refined answer: new_choice + explanation
-C.3.4 Chain-of-Thought Kojima et al. (2022)
-The CoT prompt template is provided below.
+C.5.4 Chain-of-Thought Kojima et al. (2022)
 A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}.
@@
 Explanations:...
-C.3.5 Chain-of-Opinion (Ours)
-Our CoO prompt template is provided below.
-A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}.
-Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step by explaining and analyzing each of the Opinions and Demographic Information: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A.
-or B.
-or C.
-or D.
-or E....
-Explanations:...
-C.4 Prompting Costs for API Models
-Our prompting costs for API models are reported in table 10.
-We observe that for ChatGPT and ChatGPT-Instruct, ChOiRe costs around $7$ and $10$ more US\$ dollars in total compared to baseline DIO-top8 + SC.
-However, these extra amounts of costs are worth it because we gain significant improvements over all the baselines and especially huge improvements for some topics.
-Additionally, for GPT-4, ChOiRe costs a similar price with the baseline DIO-top8 while DIO-top8 + SC costs nearly double our price.
-This is because we perform the FEA and LLMtop-$K$ steps of ChOiRe by ChatGPT, which are relatively cheap.
-DIO-top8 DIO-top8 + CoT DIO-top8 + SC ChOiRe Model
-Ave. consumed #tokens 562.72 623.62 995.89 3142.86 ChatGPT
-Total US\$ 3.01 3.73 6.82 13.95 ChatGPT
-Ave. consumed #tokens 562.72 630.58 1019.31 3121.72 ChatGPT-Instruct
-Total US\$ 3.12 3.84 7.11 19.99 ChatGPT-Instruct
-Ave. consumed #tokens 559.27 - 1021.14* 3180.82 GPT-4
-Total US\$ 91.19 - 226.15* 123.30 GPT-4
-Table 10: Prompting cost analysis of ChOiRe and other baselines as of 1st Feb 2024. * denotes our estimation on $50$ random samples.
 Appendix D Human Evaluation
-D.1 Human Rating Instructions
-Our details of human rating instructions are provided in table 11 for all the criteria.
+D.1 Human Grading Instructions
+Our details of human rating instructions are provided in Tab.
+12 for all the criteria.
 It is worth noting that selecting all features can’t get a high FEA Satisfaction score, according to our instructions.
@@
 3: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question is less than or equal to 1.
-1: The CoO has limited or flawed reasoning thoughts with inadequate support.
-CoO Reasonableness 2: The CoO has some reasoning thoughts with decent support but room for improvement.
-3: The CoO has strong, clear, and well-supported reasoning thoughts with a comprehensive understanding.
-1: The generated CoO explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
-CoO Follow the Instruction 2: The generated CoO explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
-3: The generated CoO explanation follows perfectly the instruction via explaining all the explicit and implicit attributes one by one.
+1: The VBN has limited or flawed values, beliefs, norms thoughts with inadequate support.
+VBN Reasonableness 2: The VBN has some values, beliefs, norms thoughts with decent support but room for improvement.
+3: The VBN has strong, clear, and well-supported values, beliefs, norms thoughts with a comprehensive understanding.
+1: The generated VBN explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
+VBN FI 2: The generated VBN explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
+3: The generated VBN explanation follows perfectly the instructions via explaining all the explicit and implicit attributes one by one.
 Table 12: Human rating instructions. FEA, LLMtop-$K$, and CoO stand for Filtering Explicit Personae Attributes, Implicit Personae Opinions Ranking, and Chain-of-Opinion reasoning (Section 4).
-Appendix E Examples
-E.1 FEA Example with ChatGPT
+Appendix E Error Analyses and Examples
+E.1 Error Analyses
+FEA misses key explicit personae.
+Despite showing promising results with FEA, LLMs sometimes misselect relevant personae.
+One such example is the top-left of Section E.2.
+We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
+LLMtop-$K$ opinions consist of less relevant ones.
+We observe LLMs frequently include less relevant, or even irrelevant opinions to the ranked list such as in Section E.2-bottom.
+We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
+LLMs may not follow the instructions to perform VBN reasoning.
+Although ChatGPT and ChatGPT-it demonstrate a robust ability to perform VBN reasoning (Tab.
+3), the same level of proficiency is not observed in Mistral and GPT-4, as exampled in Section E.2-top-right.
+We posit this disparity arises from the fact that ChatGPT and ChatGPT-it excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.
+E.2 Examples
+FEA example with ChatGPT.
 Fig.
@@
 Figure 5: FEA example with ChatGPT.
-E.2 Example of High Disagreement between Rankings
+Example of high disagreement between rankings.
 Fig.
@@
 Figure 6: Example of the high disagreement between orders by semantic similarity scores and LLM (ChatGPT).
-E.3 Example of Inconsistent Answers Generated by CoT
+Example of inconsistent answers generated by CoT.
 Fig.
@@
 Figure 7: Example of the inconsistent answers generated by ChatGPT with Chain-of-Thought.
-E.4 Example of Chain of Opinion Reasoning
-Fig. 11 presents an example of the answer generated by ChatGPT using Chain of Opinion (ours) versus Chain of Thought Wei et al. (2022) prompting methods.
-Figure 11: Example of an answer generated by Chain of Opinion versus Chain of Thought prompting with ChatGPT.
-E.5 Example of Answer Consistency with Dynamic Numbers of Opinions
-Fig. 12 shows an example of the answer generated by GPT-4 using Chain of Opinion (ours) reasoning with different numbers of provided historical opinions.
-Figure 12: Example of our answer consistency technique (ours), generated by GPT-4.
-E.6 Error Analysis Examples
+Human evaluation example.
 Fig.
-13 illustrates our error analysis examples of ChOiRe with ChatGPT.
+8 illustrates our human evaluation example of COO with ChatGPT.
 The top-left frame is an example of FEA missing key explicit personae.
@@
 The top-right rectangular is an example from GPT-4, showing that it does not follow human instructions to predict opinion via chain-of-opinion reasoning.
-Figure 13: Error analysis examples of ChOiRe with ChatGPT.
+Figure 8: Error analysis examples of COO with ChatGPT.